### v1

# Variable Substitution and Bilinear Programming for Aligning Partially Overlapping Point Sets

Abstract
In many applications, the demand arises for algorithms capable of aligning partially overlapping point sets while remaining invariant to the corresponding transformations. This research presents a method designed to meet such requirements through minimization of the objective function of the robust point matching (RPM) algorithm. First, we show that the RPM objective is a cubic polynomial. Then, through variable substitution, we transform the RPM objective to a quadratic function. Leveraging the convex envelope of bilinear monomials, we proceed to relax the resulting objective function, thus obtaining a lower bound problem that can be conveniently decomposed into distinct linear assignment and low-dimensional convex quadratic program components, both amenable to efficient optimization. Furthermore, a branch-and-bound (BnB) algorithm is devised, which solely branches over the transformation parameters, thereby boosting convergence rate. Empirical evaluations demonstrate better robustness of the proposed methodology against non-rigid deformation, positional noise, and outliers, particularly in scenarios where outliers remain distinct from inliers, when compared with prevailing state-of-the-art approaches.
keywords:
branch-and-bound, partial overlap , bilinear monomial , point set registration, convex envelope, linear assignment
journal: Neurocomputing
1 Introduction
Aligning 2D/3D point sets is crucial across computer vision, robotics, and remote sensing applications. Apart from disturbances like non-rigid deformation and positional noise, this task becomes especially challenging when the two point sets only partial overlap and lack initial coarse alignment.
A practical approach to addressing the point set registration challenge involves minimizing the RPM objective [1]. This objective is characterized by a mixed linear assignment-least squares formulation:
$E(P,\boldsymbol{\theta})=\sum_{i,j}p_{ij}\|\mathbf{y}_{j}-\mathbf{T}(\mathbf{x}_{i}|\boldsymbol{\theta})\|^{2}$
Here, ${\mathbf{x}_{i}}$ and ${\mathbf{y}_{j}}$ represent two point sets intended for alignment, $p_{ij}\in{0,1}$ indicates point correspondence, and $\mathbf{T}$ denotes a spatial transformation parameterized by $\boldsymbol{\theta}$. Lian et al. demonstrated that in scenarios where one point set can be embedded within another, after spatial transformations are eliminated, the RPM objective can be reduced to a concave quadratic function of point correspondence, exhibiting a low-rank structure [2]. This simplification facilitates the implementation of efficient branch-and-bound (BnB) based global optimization strategies.
The algorithm introduced in [2] is extended to handle partial overlap in [3]. Here, the objective function takes the form of a non-quadratic concave function of point correspondence variable, exhibiting a low-rank structure. This structure enables application of efficient BnB optimization techniques. However, the algorithm’s branching space is a projected space of point correspondence, characterized by a dimensionality surpassing that of the transformation parameters. This results in a slower convergence rate for the method. Additionally, to ensure concavity of the objective function over the initial bounding region, the method requires the regularization of transformations, which in turn demands prior knowledge of transformation parameters. Consequently, the method lacks invariance to corresponding transformations, presenting a notable drawback.
[…]
Figure 1: Derivation of a lower bound function of the RPM objective function via variable substitution and bilinear relaxation.
In response to this challenge, Lian and Zhang utilized the inherent constraints of rigid transformations to devise a novel objective function for the point correspondence variable. This function exhibits concavity across the entire solution space, eliminating the need for regularization of transformations [4]. Alternatively, Lian et al. employed the polyhedral annexation (PA) algorithm to minimize the RPM objective [5]. PA offers an advantage over BnB by operating exclusively within the concavity region of the objective, thus eliminating the need for transformation regularization. However, both approaches [4] and [5] still face the challenge inherited from [3] of branching within a high-dimensional projected space of point correspondence, resulting in sluggish convergence rates.
The methods mentioned above all share a common practice: they remove spatial transformations from the RPM objective, resulting in a function based solely on point correspondence variable. However, this simplification leads to a complex function, posing challenges such as high-dimensional branching spaces and intricate optimization strategies. An alternative approach, as demonstrated by [6], suggests retaining spatial transformations. In this method, trilinear and bilinear relaxation are employed to derive a lower bound function. This approach offers several advantages, including its ability to handle partial overlap, its invariance to corresponding transformations, and its utilization of a low-dimensional branching space.
Following the approach outlined in [6], we introduce a new Branch-and-Bound (BnB) algorithm tailored for optimizing the RPM objective function. This algorithm boasts the following features:
  • The algorithm proposed by [6] is complex because it necessitates reformulating the RPM objective function to ensure that one of the variables within the trilinear monomial includes the origin within its range. In contrast, as illustrated in Fig. 1, our method exclusively utilizes bilinear relaxation to establish the lower bound function, eliminating the necessity to fulfill this condition. This results in a streamlined algorithm compared to [6].
  • The optimization of the lower bound function can be broken down into two distinct tasks: linear assignment and low-dimensional convex quadratic optimization related to spatial transformation variables. Both of these tasks can be efficiently solved with global optimality assurances.
  • Our approach inherits the advantageous characteristics outlined in [6], including its ability in handling partial overlap, invariance to corresponding transformations, and utilization of a low-dimensional branching space.
The paper is structured as follows: We begin with a review of related work (Sec. 2). Following that, we introduce the theoretical framework necessary for developing our lower bound functions (Sec. 3). Next, we delve into our objective function (Sec. 4 and optimization strategy (5). Subsequently, we apply the proposed algorithm to 2D and 3D registration (Sec. 6 and Sec. 7). Finally, we present our experimental results (Sec. 8) and conclude the paper (Sec. 9).
2 Related work
A comprehensive survey of point set registration methods is available in [7]. Here, we focus solely on works closely related to ours.
2.1 Local methods
Simultaneous pose and correspondence. The widely known ICP method [8] iterates between recovering point correspondence and updating spatial transformation. Despite its efficiency, the discrete nature of point correspondence often traps the method in local minima. Zhang et al. enhanced the robustness of ICP by constraining the rotation angle at each iterative step [9]. Liang introduced RGB information to the registration process of ICP [10]. The RPM method [1] relaxes point correspondence to be fuzzily valued and employs deterministic annealing (DA) to gradually recover point correspondence. Instead of DA, Sofka et al. utilized the covariance matrix of the transformation parameters to determine the fuzziness of point correspondence, resulting in better robustness to missing or extraneous structures [11]. Ma et al. employed L2E to robustly estimate transformation for non-rigid registration [12]. Our method is related to the aforementioned methods by also modeling spatial transformation and correspondences. However, our objective function is globally minimized, which enhances our method’s robustness to disturbances.
Correspondence-free methods. The CPD method [13] frames point matching as fitting a Gaussian mixture model (GMM) representing one point set to another. GMMREG [14] utilizes two GMMs to represent two point sets and minimizes the $L_{2}$-norm distance between them. The SVGMs method represents point sets using sparse Gaussian components [15]. Efficiency in GMM-based methods is enhanced by employing filtering to solve the correspondence problem [16]. The density variation problem of point sets is addressed in [17] by modeling scene structure. Hierarchical Gaussian mixture representation improves speed and accuracy of registration [18].
2.2 Global methods
BnB based methods. The Branch-and-Bound (BnB) algorithm stands as a widely-used global optimization technique. Rooted in Lipschitz theory, BnB is applied to align 3D shapes [19], albeit with limitations regarding outliers or occlusions. By capitalizing on the structure of 3D rigid motions geometry, Go-ICP [20] globally optimizes the objective of ICP. GOGMA [21] achieves point set registration by aligning Gaussian Mixture Models (GMMs) constructed from the original point sets. Employing stereographic projection, the fast rotation search (FRS) method efficiently recovers 3D rotation between two point sets [22]. Bayesian nonparametric mixture and a novel approach of tessellating rotation space are introduced in [23], leading to an efficient alignment algorithm. Our BnB-based method distinguishes itself from the aforementioned methods by also incorporating point correspondences. This feature empowers our method to better handle non-rigid deformation.
Mismatch removal methods. An alternative avenue of research focuses on recovering spatial transformations given putative point correspondences. The fast global registration (FGR) method [24] optimizes a robust objective grounded in the duality between line processes and robust estimation. GORE [25] eliminates a large portion of outliers using geometric operations before invoking RANSAC. TEASER++ [26] employs a graph-theoretic approach to decouple rotation, scale, and translation estimation. It adopts a truncated least squares formulation for each subproblem, enabling effective mismatch removal.
2.3 Deep learning methods
Correspondence-free methods. The Lucas–Kanade algorithm aligns global features of point sets generated by a PointNet network in [27]. This approach is refined in [28] by employing a decoder to convert the generated global features back into point sets. Ensuring fidelity, Chamfer distances between the decoded point sets and the original ones are minimized.
Correspondence-based methods. The FCGF method [29] employs sparse high-dimensional convolutions to extract dense feature descriptors from point clouds. This method significantly outperforms many hand-crafted features and PointNet-based methods. Additionally, it surpasses a 3D convolution-based method [30]. The deep global registration method [31] utilizes FCGF feature descriptors for point cloud registration. DCP [32] utilizes DGCNN [33] and Transformer [34] to extract features for each point set, with SVD employed to recover rigid transformation. PRNet [35] extends DCP to handle partially overlapping point sets iteratively. IDAM [36] considers both geometric and distance features during its point matching process. RIENet [37] calculates feature-to-feature correspondences with neighborhood consensus. Finally, [38] performs partial point set registration by generating matching credibility on the global semantic level and learning geometric features on the local structural level.
3 Bilinear convex envelope linearization
The convex envelope of a bilinear monomial $xy$ subject to constraints $\underline{x}\leq x\leq\overline{x}$ and $\underline{y}\leq y\leq\overline{y}$ can be described as:
$\begin{cases}(xy)_{l1}\triangleq\underline{x}y+x\underline{y}-\underline{x}\underline{y}\leq xy,\\ (xy)_{l2}\triangleq\overline{x}y+x\overline{y}-\overline{x}\overline{y}\leq xy\end{cases}$ (1)
There are two ways of utilizing this formula when devising the lower bound problem: (1) Integration of the term $\max\{(xy)_{l1},(xy)_{l2}\}$ into the lower bound objective function. Nevertheless, this incorporation can lead to non-linearity of the resultant function. (2) Formulation of a subproblem in the format $\min\{\alpha|\alpha\geq(xy)_{l1},\alpha\geq(xy)_{l2}\}$ within the lower bound problem. However, this approach increases the complexity of constraints and raises concerns regarding the feasibility of the resulting problem for efficient optimization via combinatorial optimization algorithms.
Hence, in alignment with the methodology proposed in [6], we advocate for using the average of the convex envelope components, as delineated in (2):
$(xy)_{l}\triangleq\frac{1}{2}\sum_{i}(xy)_{li}\leq xy$ (2)
as the lower bound function for $xy$. This choice has the advantage that the lower bound function is linear, thereby facilitating efficient optimization by combinatorial optimization algorithms.
[6] also shows that for $(xy)_{l}$ to converge to $xy$, we only need to branch over one variable $x$ or $y$ that makes up the bilinear monomial $xy$. This is a crucial property and the good convergence of the proposed algorithm hings on this fact.
4 Problem formulation
Suppose there are two point sets $\mathscr{X}=\{\mathbf{x}_{i},1\leq i\leq n_{x}\}$ and $\mathscr{Y}=\{\mathbf{y}_{j},1\leq j\leq n_{y}\}$ in $\mathbb{R}^{n_{d}}$ to be aligned, where the column vectors $\mathbf{x}_{i}$ and $\mathbf{y}_{j}$ represent the coordinates of points $i$ and $j$, repectively. Based on [3], this problem can be modeled as a mixed linear assignment--least square problem:
$\min E(\mathbf{P},\boldsymbol{\theta})=\sum_{i,j}p_{ij}\|\mathbf{y}_{j}-\mathbf{J}(\mathbf{x}_{i})\boldsymbol{\theta}\|^{2}$ (3a)
$={\boldsymbol{\theta}}^{\top}\mathbf{J}^{\top}(\text{diag}(\mathbf{P}\mathbf{1}_{n_{y}})\otimes\mathbf{I}_{d})\mathbf{J}{\boldsymbol{\theta}}-2{\boldsymbol{\theta}}^{\top}\mathbf{J}^{\top}(\mathbf{P}\otimes\mathbf{I}_{d})\mathbf{y}+\mathbf{1}_{n_{x}}^{\top}\mathbf{P}\widetilde{\mathbf{y}}$ (3b)
[…]
where
$\boldsymbol{\theta}$ : parameters of the spatial transformation $\mathbf{T}(\mathbf{x}_{i}|\boldsymbol{\theta})=\mathbf{J}(\mathbf{x}_{i})\boldsymbol{\theta}$. $\mathbf{J}(\mathbf{x}_{i})$ : the Jacobian matrix at point $\mathbf{x}_{i}$. $\mathbf{I}_{d}$: the $d\times d$ identity matrix. $\mathbf{1}_{n_{x}}$: the $n_{x}$-dimensional vector of all ones. $\otimes$: the Kronecker product. $\mathbf{P}=\{p_{ij}\}$ : assignment matrix, with element $p_{ij}{=}1$ or 00 indicating that $\mathbf{x}_{i}$ matches $\mathbf{y}_{j}$ or not. $\mathbf{1}_{n_{x}}^{\top}\mathbf{P}\mathbf{1}_{n_{y}}{=}n_{p}$: the number of matches is a priori known to be a constant integer $n_{p}$. $\underline{\boldsymbol{\theta}}$ (resp. $\overline{\boldsymbol{\theta}}$) : the lower (resp. upper) bound of $\boldsymbol{\theta}$.
$\mathbf{y}\triangleq\begin{bmatrix}\mathbf{y}_{1}^{\top},\dots,\mathbf{y}_{n_{y}}^{\top}\end{bmatrix}^{\top}$: a vector. $\widetilde{\mathbf{y}}\triangleq\begin{bmatrix}\|\mathbf{y}_{1}\|_{2}^{2},\dots,\|\mathbf{y}_{n_{y}}\|_{2}^{2}\end{bmatrix}^{\top}$: a vector. $\mathbf{J}\triangleq\begin{bmatrix}\mathbf{J}^{\top}(\mathbf{x}_{1}),\dots,\mathbf{J}^{\top}(\mathbf{x}_{n_{x}})\end{bmatrix}^{\top}$: a matrix.
For ease of derivation in the sequel, we need to convert matrix $\mathbf{P}$ into a vector $\mathbf{p}$. Accordingly, the objective function takes the following form [3]:
$E(\mathbf{p},\boldsymbol{\theta})=\boldsymbol{\theta}^{\top}\text{mat}(\mathbf{B}\mathbf{p})\boldsymbol{\theta}-2\boldsymbol{\theta}^{\top}\mathbf{A}\mathbf{p}+\boldsymbol{\rho}^{\top}\mathbf{p}$ (4)
where
$\mathbf{p}\triangleq\text{vec}(\mathbf{P})$: a vector. $\boldsymbol{\rho}\triangleq\mathbf{1}_{n_{x}}\otimes\widetilde{\mathbf{y}}$: a vector. $\mathbf{A}\triangleq(\mathbf{J}^{\top}\otimes\mathbf{y}^{\top})\mathbf{W}^{n_{x},n_{y}}_{n_{d}}$: a matrix. $\mathbf{B}\triangleq(\mathbf{J}_{2}^{\top}\otimes\mathbf{I}_{n_{\theta}})\mathbf{W}^{n_{x},1}_{n_{\theta}}(\mathbf{I}_{n_{x}}\otimes\mathbf{1}_{n_{y}}^{\top})$: a matrix. $\mathbf{J}_{2}\triangleq\begin{bmatrix}\mathbf{J}(\mathbf{x}_{1})^{\top}\mathbf{J}(\mathbf{x}_{1}),\ldots,\mathbf{J}(\mathbf{x}_{n_{x}})^{\top}\mathbf{J}(\mathbf{x}_{n_{x}})\end{bmatrix}^{\top}$: a matrix. $\mathbf{W}^{m,n}_{d}\triangleq\mathbf{I}_{m}\otimes\begin{bmatrix}\mathbf{I}_{n}\otimes(\mathbf{e}_{d}^{1})^{\top},\dots,\mathbf{I}_{n}\otimes(\mathbf{e}_{d}^{d})^{\top}\end{bmatrix}^{\top}$: a matrix.
$\text{vec}(\cdot)$: vectorization of a matrix by concatenating rows. $n_{\theta}$: the dimensionality of ${\boldsymbol{\theta}}$. $\mathbf{e}_{d}^{i}$: the $d$-dimensional column vector with only one nonzero unit element at the $i$-th position. $\text{mat}(\cdot)$: reconstructs a matrix from a vector, which is the inverse of the operator $\text{vec}(\cdot)$.
Since $\mathbf{1}^{\top}_{n_{x}n_{y}}\mathbf{p}=n_{p}$, a constant, for rows of $\mathbf{B}$ containing identical elements, the result of them multiplied by $\mathbf{p}$ can be replaced by constants. Also, redundant rows can be removed. Since $\text{mat}(\mathbf{B}\mathbf{p})$ is a symmetric matrix, $\mathbf{B}$ surely contains redundant rows. Based on the above analysis, we therefore denote $\mathbf{B}_{2}$ as the matrix formed as a result of $\mathbf{B}$ removing such rows. (Please refer to Sec. 6 and 7 for examples of $\mathbf{B}_{2}$). Consequently, $E$ can be rewritten as [6]
$E(\mathbf{p},\boldsymbol{\theta})=\boldsymbol{\theta}^{\top}\left[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})+\mathbf{C}\right]\boldsymbol{\theta}-2\boldsymbol{\theta}^{\top}\mathbf{A}\mathbf{p}+\boldsymbol{\rho}^{\top}\mathbf{p}$ (5)
where the nonzero elements of the constant matrix $\mathbf{C}$ correspond to the rows of $\mathbf{B}$ containing identical elements. The elements of the constant matrix $\mathbf{K}$ take values in $\{0,1\}$ and record the correspondences between the rows of $\mathbf{B}_{2}$ and those of $\mathbf{B}$.
[…]
Figure 2: Derivation of a lower bound function $E_{l}(\mathbf{p},\boldsymbol{\theta})$ of $E(\mathbf{p},\boldsymbol{\theta})$ via variable substitution and bilinear relaxation.
5 Optimization
In this section, we begin by applying the bilinear relaxation formula, introduced in Section 3, to derive a lower bound function for our objective. This lower bound function’s optimization can then be divided into separate optimizations over $\boldsymbol{\theta}$ and $\mathbf{p}$ (discussed in Section 5.1). Subsequently, we demonstrate that it suffices to branch solely over $\boldsymbol{\theta}$ (explained in Section 5.2) and introduce a more effective method for computing the upper bound (outlined in Section 5.4). Lastly, we introduce the BnB algorithm (presented in Section 5.5), which offers global optimization for the objective function.
5.1 Lower bound function derivation
We derive a lower bound function of $E(\mathbf{p},\boldsymbol{\theta})$ via variable substitution and bilinear relaxation, as illustrated in Fig. 2. In the following, we explain each step of the approach in detail.
By introducing new variables $\boldsymbol{\Theta}\triangleq\boldsymbol{\theta}\boldsymbol{\theta}^{\top}$, $\boldsymbol{\Gamma}\triangleq\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})$ and $\boldsymbol{\eta}\triangleq-2\mathbf{A}\mathbf{p}$, we can rewrite the above objective function as a quadratic function:
$E(\mathbf{p},\boldsymbol{\theta},\boldsymbol{\Theta},\boldsymbol{\Gamma},\boldsymbol{\eta})=\text{trace}(\boldsymbol{\Gamma}\boldsymbol{\Theta})+\boldsymbol{\theta}^{\top}\mathbf{C}\boldsymbol{\theta}+\boldsymbol{\theta}^{\top}\boldsymbol{\eta}{+}\boldsymbol{\rho}^{\top}\mathbf{p}$ (6)
We note that $E$ contains bilinear terms $\text{trace}(\boldsymbol{\Gamma}\boldsymbol{\Theta})$ and $\boldsymbol{\theta}^{\top}\boldsymbol{\eta}$, which interconnect $\boldsymbol{\Gamma}$ and $\boldsymbol{\Theta}$ (resp. $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$), making optimization challenging. To tackle this challenge, we use the bilinear relaxation formula introduced in Section 3 to relax these terms into linear ones. The first relaxation is:
${{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boxed{\text{trace}(\boldsymbol{\Gamma}\boldsymbol{\Theta})}}}$${{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boxed{\text{trace}(\mathbf{H}^{0}\boldsymbol{\Theta})+\text{trace}(\mathbf{H}^{1}\boldsymbol{\Gamma})+\text{trace}(\mathbf{H}^{2}\mathbf{1}\mathbf{1}^{\top})}}}$bilinearrelaxation (7)
where in accordance with equations (1) and (2), the coefficients can be calculated as
$\mathbf{H}^{0}=\frac{\underline{\boldsymbol{\Gamma}}+\overline{\boldsymbol{\Gamma}}}{2},\mathbf{H}^{1}=\frac{\underline{\boldsymbol{\Theta}}+\overline{\boldsymbol{\Theta}}}{2},\mathbf{H}^{2}=-\frac{\underline{\boldsymbol{\Gamma}}\underline{\boldsymbol{\Theta}}+\overline{\boldsymbol{\Gamma}}\overline{\boldsymbol{\Theta}}}{2}$ (8)
Here we use underline (resp. overline) to denote the lower (resp. upper) bounds of matrices $\boldsymbol{\Theta}$ and $\boldsymbol{\Gamma}$. It is also worth noting that the range $[\underline{\boldsymbol{\Theta}}$, $\overline{\boldsymbol{\Theta}}]$ of $\boldsymbol{\Theta}$ can be derived from the range $[\underline{\boldsymbol{\theta}},\overline{\boldsymbol{\theta}}]$ of $\boldsymbol{\theta}$ based on relation $\boldsymbol{\Theta}=\boldsymbol{\theta}\boldsymbol{\theta}^{\top}$, as explained in Section 5.3.
The second relaxation is:
${{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boxed{\boldsymbol{\theta}^{\top}\boldsymbol{\eta}}}}$${{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boxed{\begin{matrix}(\mathbf{g}^{0})^{\top}\boldsymbol{\theta}+(\mathbf{g}^{1})^{\top}\boldsymbol{\eta}+\mathbf{1}_{n_{\theta}}^{\top}\mathbf{g}^{2}\end{matrix}}}}$bilinearrelaxation (9)
where in accordance with equations (1) and (2), the coefficients can be calculated as
$\mathbf{g}^{0}=\frac{\underline{\boldsymbol{\eta}}+\overline{\boldsymbol{\eta}}}{2},\mathbf{g}^{1}=\frac{\underline{\boldsymbol{\theta}}+\overline{\boldsymbol{\theta}}}{2},\mathbf{g}^{2}=\frac{\underline{\boldsymbol{\eta}}\underline{\boldsymbol{\theta}}+\overline{\boldsymbol{\eta}}\overline{\boldsymbol{\theta}}}{2}$ (10)
Here, the underline and overline have the same meaning as before.
Taking into account the above relaxations, we get the following lower bound function:
$E_{l}(\mathbf{p},\boldsymbol{\theta},\boldsymbol{\Theta},\boldsymbol{\Gamma},\boldsymbol{\eta})=$ $\text{trace}(\mathbf{H}^{0}\boldsymbol{\Theta})+\text{trace}(\mathbf{H}^{1}\boldsymbol{\Gamma})+\text{trace}(\mathbf{H}^{2}\mathbf{1}\mathbf{1}^{\top})$
$+\boldsymbol{\theta}^{\top}\mathbf{C}\boldsymbol{\theta}+\boldsymbol{\theta}^{\top}\mathbf{g}^{0}+(\mathbf{g}^{1})^{\top}\boldsymbol{\eta}+\mathbf{1}^{\top}\mathbf{g}^{2}{+}\boldsymbol{\rho}^{\top}\mathbf{p}$ (11)
By substituting back $\boldsymbol{\theta}\boldsymbol{\theta}^{\top}\Leftarrow\boldsymbol{\Theta}$, $\text{mat}(K\mathbf{B}_{2}\mathbf{p})\Leftarrow\boldsymbol{\Gamma}$ and $-2\mathbf{A}\mathbf{p}\Leftarrow\boldsymbol{\eta}$, we get the final form of the lower bound function:
$E_{l}(\mathbf{p},\boldsymbol{\theta})=$ $\boldsymbol{\theta}^{\top}\mathbf{H}^{0}\boldsymbol{\theta}+\text{trace}(\mathbf{H}^{1}\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p}))+\text{trace}(\mathbf{H}^{2}\mathbf{1}\mathbf{1}^{\top})$
$+\boldsymbol{\theta}^{\top}\mathbf{C}\boldsymbol{\theta}+\boldsymbol{\theta}^{\top}\mathbf{g}^{0}+(\mathbf{g}^{1})^{\top}(-2\mathbf{A}\mathbf{p})+\mathbf{1}_{n_{\theta}}^{\top}\mathbf{g}^{2}+\boldsymbol{\rho}^{\top}\mathbf{p}$ (12)
It is apparent that the minimization of $E_{l}$ under constraints (3c) can be decomposed into separate optimizations over $\mathbf{p}$ and $\boldsymbol{\theta}$.
The optimization over $\mathbf{p}$ is
$\min_{p}\ \text{trace}(\mathbf{H}^{1}\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p}))+(\mathbf{g}^{1})^{\top}(-2\mathbf{A}\mathbf{p})+\boldsymbol{\rho}^{\top}\mathbf{p}$ (13a)
$s.t.\ \mathbf{p}\in\Omega$ (13b)
where $\Omega$ denotes the feasible region of $\mathbf{p}$ as defined in (3c). This is an $n_{p}-$cardinality linear assignment problem and can be transformed into a standard linear assignment problem [39] and then efficiently solved by combinatorial optimization algorithms such as LAPJV [40].
The optimization over $\boldsymbol{\theta}$ is:
$\min_{\boldsymbol{\theta}}\ \boldsymbol{\theta}^{\top}(\mathbf{H}^{0}+\mathbf{C})\boldsymbol{\theta}+\boldsymbol{\theta}^{\top}\mathbf{g}^{0}$ (14a)
$s.t.\ \underline{\boldsymbol{\theta}}\leq\boldsymbol{\theta}\leq\overline{\boldsymbol{\theta}}$ (14b)
In the next section, we will delve into the solution for this problem.
5.2 Branching strategy
According to the result stated in Section 3, for the bilinear term $\boldsymbol{\theta}^{\top}\boldsymbol{\eta}$, for its linear lower bound function to converge to it, it suffices to branch over $\boldsymbol{\theta}$ and leave the range of $\boldsymbol{\eta}$ fixed. Here, utilizing $\boldsymbol{\eta}=-2\mathbf{A}\mathbf{p}$, the fixed range $[\underline{\boldsymbol{\eta}},\overline{\boldsymbol{\eta}}]$ of $\boldsymbol{\eta}$ can be computed as
$\underline{\eta}_{i}=\min_{\mathbf{p}\in\Omega}(-2\mathbf{A}\mathbf{p})_{i}$ (15)
$\overline{\eta}_{i}=\max_{\mathbf{p}\in\Omega}(-2\mathbf{A}\mathbf{p})_{i}$ (16)
Likewise, for the bilinear term $\text{trace}(\boldsymbol{\Gamma}\boldsymbol{\Theta})$, for its linear lower bound function to converge to it, we only need to branch over $\boldsymbol{\theta}$ (since the range of $\boldsymbol{\theta}$ determines the range of $\boldsymbol{\Theta}$ as explained in Sec. 5.3) and leave the range of $\boldsymbol{\Gamma}$ fixed. Here, utilizing $\boldsymbol{\Gamma}=\text{mat}(K\mathbf{B}_{2}\mathbf{p})$, the fixed range $[\underline{\boldsymbol{\Gamma}},\overline{\boldsymbol{\Gamma}}]$ of $\boldsymbol{\Gamma}$ can be computed as follows:
$\underline{\Gamma}_{ij}=\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$ (17)
$\overline{\Gamma}_{ij}=\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$ (18)
With thess choice of $\underline{\boldsymbol{\Gamma}}$ and $\overline{\boldsymbol{\Gamma}}$, it has been empirically demonstrated that
$\mathbf{H}^{0}+\mathbf{C}=\frac{\left(\underline{\boldsymbol{\Gamma}}+\mathbf{C}\right)+\left(\overline{\boldsymbol{\Gamma}}+\mathbf{C}\right)}{2}\succeq 0$ (19)
Please refer to Section 5.6 for relevant experiments and Appendix for a weaker assertion. This result implies that problem (14a) (14b) is now a convex quadratic program. Consequently, utilizing solvers for convex quadratic programs, such as MATLAB’s quadprog function, guarantees global optimality. We will address the proof of (19) in our future research.
We also note that (15), (16), (17) and (18) are $n_{p}-$cardinlaity linear assignment problems, which can be efficiently solved by the aforementioned algorithms.
In conclusion, in our BnB algorithm, we only need to branch over $\boldsymbol{\theta}$. Therefore, the dimension of the branching space of our BnB algorithm is low and the proposed algorithm can converge quickly.
5.3 The range of $\Theta$ from the range of $\theta$
Given the relationship $\boldsymbol{\Theta}=\boldsymbol{\theta}\boldsymbol{\theta}^{\top}$, and the specified range $\left[\underline{\boldsymbol{\theta}},\overline{\boldsymbol{\theta}}\right]$ of $\boldsymbol{\theta}$, we can determine the range $\left[\underline{\boldsymbol{\Theta}},\overline{\boldsymbol{\Theta}}\right]$ of $\boldsymbol{\Theta}$ as follows:
$[\underline{\Theta_{ij}},\overline{\Theta_{ij}}]=\begin{cases}[\underline{\theta_{i}},\overline{\theta_{i}}]\cdot[\underline{\theta_{j}},\overline{\theta_{j}}],&\text{if}\ i\neq j\\ [\underline{\theta_{i}},\overline{\theta_{i}}]^{2},&\text{if}\ i=j\end{cases}$
Here, employing interval arithmetic, the product operator is defined as:
$[\underline{\theta_{i}},\overline{\theta_{i}}]\cdot[\underline{\theta_{j}},\overline{\theta_{j}}]=[\min(\underline{\theta_{i}}\underline{\theta_{j}},\underline{\theta_{i}}\overline{\theta_{j}},\overline{\theta_{i}}\underline{\theta_{j}},\overline{\theta_{i}}\overline{\theta_{j}}),\max(\underline{\theta_{i}}\underline{\theta_{j}},\underline{\theta_{i}}\overline{\theta_{j}},\overline{\theta_{i}}\underline{\theta_{j}},\overline{\theta_{i}}\overline{\theta_{j}})]$
Additionally, the square operator is defined as:
$[\underline{\theta_{i}},\overline{\theta_{i}}]^{2}=\begin{cases}[\min(\underline{\theta_{i}}^{2},\overline{\theta_{i}}^{2}),\max(\underline{\theta_{i}}^{2},\overline{\theta_{i}}^{2})],&\text{if}\ 0\notin[\underline{\theta_{i}},\overline{\theta_{i}}]\\ [0,\max(\underline{\theta_{i}}^{2},\overline{\theta_{i}}^{2})],&\text{if}\ 0\in[\underline{\theta_{i}},\overline{\theta_{i}}]\end{cases}$
5.4 Upper bound
When we substitute the solutions for $\mathbf{p}$ and $\boldsymbol{\theta}$ obtained during the computation of the lower bound into $E(\mathbf{p},\boldsymbol{\theta})$, we can derive an upper bound. However, the solution for $\boldsymbol{\theta}$ obtained in this process might significantly differ from the optimal $\boldsymbol{\theta}$, especially in the early iterations of the BnB algorithm. This discrepancy can lead to a poor computed upper bound.
To address this issue, we can rewrite the objective function $E(\mathbf{p},\boldsymbol{\theta})$ as a function of $\mathbf{p}$ alone by eliminating $\boldsymbol{\theta}$, based on a result from [6]:
$E(\mathbf{p})=-\mathbf{p}^{\top}\mathbf{A}^{\top}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})+\mathbf{C}]^{-1}\mathbf{A}\mathbf{p}+\boldsymbol{\rho}^{\top}\mathbf{p}.$ (20)
Therefore, by substituting the computed $\mathbf{p}$ into this function, we obtain a value for $E$, which serves as an upper bound.
5.5 Branch-and-Bound
Based on the aforementioned preparations, we are now poised to apply the BnB algorithm to optimize $E$. To initiate the process, we utilize the initial range of $\boldsymbol{\theta}$ to form the initial hypercube $M$. Subsequently, during each iteration of the algorithm, we identify the hypercube that yields the lowest lower bound among all available hypercubes. This selected hypercube is then further subdivided to enhance the global lower bound of the problem. Simultaneously, the upper bound is refined by assessing $E(\mathbf{p})$ using solutions generated during the computation of the lower bounds. For a comprehensive overview, refer to Algorithm 1.
1
2Set tolerance error $\epsilon>0$ and initial hypercube $M$. Let $\mathscr{M}_{1}=\mathscr{N}_{1}=\{M\}$ where $\mathscr{M}_{k}$ and $\mathscr{N}_{k}$ denote the collection of all hypercubes and the collection of active hypercubes at iteration $k$, respectively.
3for $k=1,2,\ldots$ do
4
5      For each hypercube $M\in\mathscr{N}_{k}$, minimize the lower bound function $E_{l}$ to obtain the optimal point correspondence solution $\mathbf{p}(M)$ and the optimal value $\beta(M)$. $\beta(M)$ is a lower bound for $M$.
6      Let $\mathbf{p}^{k}$ be the best among all feasible solutions so far encountered: $\mathbf{p}^{k-1}$ and all $\mathbf{p}(M)$ for $M\in\mathscr{N}_{k}$. Delete all $M\in\mathscr{M}_{k}$ such that $\beta(M)\geq E(\mathbf{p}^{k})-\epsilon$. Let $\mathscr{R}_{k}$ be the remaining collection of hypercubes.
7      If $\mathscr{R}_{k}=\emptyset$, terminate: $\mathbf{p}^{k}$ is the global $\epsilon-$minimum solution.
8      Select the hypercube $M$ yielding the lowest lower bound and divide it into two sub-hypercubes $M_{1}$, $M_{2}$ by bisecting the longest edge.
9      Let $\mathscr{N}_{k+1}=\{M_{1},M_{2}\}$ and $\mathscr{M}_{k+1}=(\mathscr{R}_{k}\backslash M)\cup\mathscr{N}_{k+1}$.
10 end for
11
Algorithm 1 A BnB algorithm for minimizing $E$
5.6 Positive semidefiniteness of $\mathbf{H}^{0}+\mathbf{C}$
To confirm the assertion that $\mathbf{H}^{0}+\mathbf{C}\succeq 0$, we utilize the separate outliers and inliers test detailed in Section 8.1.1, with an outlier to data ratio set at 0.3. We present histograms of the smallest eigenvalues of the matrix $\mathbf{H}^{0}+\mathbf{C}$ across 100 registration instances in Figure 3. The result consistently support the claim that $\mathbf{H}^{0}+\mathbf{C}\succeq 0$, underscoring the efficacy of the proposed algorithm.
Refer to caption Refer to caption
Figure 3: Histograms depicting the smallest eigenvalue of $\mathbf{H}^{0}+\mathbf{C}$ across 100 registration instances.
5.7 Convergence of the BnB algorithm
To assess the convergence of the proposed BnB algorithm, we conducted the separate outliers and inliers test detailed in Sec. 8.1.1, with an outlier to data ratio set at $0.3$. The experimental results are depicted in Fig. 4. Key observations include: 1) A narrower initial range of $\boldsymbol{\theta}$ results in a tighter duality gap, evidenced by improved upper and lower bounds. This suggests that leveraging additional information about the range of $\boldsymbol{\theta}$ could effectively reduce the duality gap in practical applications. 2) Easier problem instances, such as the fish test compared to the character test, exhibit a tighter duality gap, thus leads to faster convergence.
Properties 2) suggests that employing a fixed duality gap threshold as the stopping criterion for all types of problems may not be a good idea. It may happen that a threshold is set too small for challenging problems, leading to prolonged execution times. Conversely, it might be set too large for simpler problems, resulting in premature termination. Therefore, following [6], we adopt the maximum branching depth as our stopping criterion.
Figure 4: Upper (first row) and lower bounds (second row) generated in each iteration of RPM-BP. Our method is tested with the $n_{p}$ value chosen as the ground truth and with varying initial ranges of $\boldsymbol{\theta}$: $[\boldsymbol{\theta}_{gt}-\Delta,\boldsymbol{\theta}_{gt}+\Delta]$, where $\boldsymbol{\theta}_{gt}$ represents the ground truth $\boldsymbol{\theta}$ solution. Here, the margin $\Delta$ takes values of $0.5$, $1$, and $1.5$ respectively.
6 Application 1: 2D similarity/affine registration
Our algorithm can handle 2D similarity and affine transformations since both of them can be written as the form of $T(\mathbf{x}|\boldsymbol{\theta})=J(\mathbf{x})\boldsymbol{\theta}$. A 2D similarity transformation has the form
$T(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&-\theta_{2}\\ \theta_{2}&\theta_{1}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2}\end{bmatrix}+\begin{bmatrix}\theta_{3}\\ \theta_{4}\end{bmatrix}=\begin{bmatrix}x^{1}&-x^{2}&1&0\\ x^{2}&x^{1}&0&1\end{bmatrix}\boldsymbol{\theta}$
where $[\theta_{3},\theta_{4}]^{\top}$ is translation and $\theta_{1}=s\cos(\phi)$ and $\theta_{2}=s\sin(\phi)$ with $s$ being scale and $\phi$ being rotation angle. It can be derived that $\mathbf{B}_{2}=\mathbf{B}([1,3,4],:)$ and $\mathbf{C}=\text{diag}([0,0,n_{p},n_{p}])$.
A 2D affine transformation has the form
$T(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&\theta_{2}\\ \theta_{3}&\theta_{4}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2}\end{bmatrix}+\begin{bmatrix}\theta_{5}\\ \theta_{6}\end{bmatrix}=\begin{bmatrix}x^{1}&x^{2}&0&0&1&0\\ 0&0&x^{1}&x^{2}&0&1\end{bmatrix}\boldsymbol{\theta}$
It can be derived that $\mathbf{B}_{2}=\mathbf{B}([1,2,5,8,11],:)$ and $\mathbf{C}=\text{diag}([0,0,0,0,n_{p},n_{p}])$.
7 Application 2: 3D rigid registration
Theoretically, our method can handle 3D affine transformation since it takes the form $T(\mathbf{x}|\boldsymbol{\theta})=J(\mathbf{x})\boldsymbol{\theta}$. A 2D affine transformation has the form
$T(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&\theta_{2}&\theta_{3}\\ \theta_{4}&\theta_{5}&\theta_{6}\\ \theta_{7}&\theta_{8}&\theta_{9}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2}\\ x^{3}\end{bmatrix}+\begin{bmatrix}\theta_{10}\\ \theta_{11}\\ \theta_{12}\end{bmatrix}$
=$\begin{bmatrix}x^{1}&x^{2}&x^{3}&0&0&0&0&0&0&1&0&0\\ 0&0&0&x^{1}&x^{2}&x^{3}&0&0&0&0&1&0\\ 0&0&0&0&0&0&x^{1}&x^{2}&x^{3}&0&0&1\end{bmatrix}\boldsymbol{\theta}$
It can be derived that $\mathbf{B}_{2}=\mathbf{B}([1,2,3,10,14,15,22,27,34],:)$ and $\mathbf{C}=\text{diag}([\underbrace{0,\ldots,0}_{9\text{ zeros}},n_{p},n_{p},n_{p}])$. Nevertheless, the 3D affine transformation has many parameters, causing our method to converge slowly. Therefore, for the 3D case, we instead focus on rigid transformation instead.
[…]
$\mathbf{R}=\mathbf{I}_{3}+\frac{[\mathbf{r}]_{\times}\sin\|\mathbf{r}\|}{\|\mathbf{r}\|}+\frac{[\mathbf{r}]_{\times}^{2}(1-\cos\|\mathbf{r}\|)}{\|\mathbf{r}\|^{2}}$ (21)
where $[\cdot]_{\times}$ denotes the skew-symmetric matrix representation:
$[\mathbf{r}]_{\times}=\begin{bmatrix}0&-r_{3}&r_{2}\\ r_{3}&0&-r_{1}\\ -r_{2}&r_{1}&0\end{bmatrix}$ (22)
where $r_{i}$ is the $i$-th element of $\mathbf{r}$.
3D rigid transformation is a special case of 3D affine transformation, whereas our method is capable of handling 3D affine transformation. Therefore, we aim to adapt our method from Section 5 to accommodate 3D rigid transformation as well. Here’s the procedure: We continue to use the algorithm outlined in Section 5 for registration, where the transformation is defined as a 3D affine transformation: $T(\mathbf{x}|\mathbf{R},\mathbf{t})=\mathbf{R}\mathbf{x}+\mathbf{t}$, with $\mathbf{R}$ representing the linear part of the affine transformation. The main differences from the algorithm in Section 5 are as follows: (1) Instead of branching over $\mathbf{R}$ and $\mathbf{t}$, we branch over $\mathbf{r}$ and $\mathbf{t}$. (2) In each iteration of the BnB algorithm, given the range of $\mathbf{r}$, we calculate the range of $\mathbf{R}$ from the range of $\mathbf{r}$ based on equation (21). (The method for efficiently computing the range of $\mathbf{R}$ is discussed in Section 7.1.)
In the above algorithm, (20) is employed to compute the upper bound, assuming that the transformation can be affine. While methods exist (see Sec. 5.3 of [6]) for computing the upper bound with a rigid transformation constraint, they often result in higher upper bounds, which slows down our method’s convergence. Hence, we opt to use (20) for computing the upper bound in this work.
7.1 The range of $\mathbf{R}$ from the range of $\mathbf{r}$
Given a range $[\underline{\mathbf{r}},\overline{\mathbf{r}}]$ for $\mathbf{r}$, the computation of the range of $\mathbf{R}$, as derived from (21), involves determining $R_{ij}$ values within the bounds specified by $\underline{\mathbf{r}}$ and $\overline{\mathbf{r}}$. This computation can be facilitated through optimization solvers such as the Matlab function $fmincon$. However, this process can be burdensome, particularly when executed iteratively as a subroutine of the Branch and Bound (BnB) algorithm. In order to alleviate this computational burden, we propose the utilization of precomputation techniques to enhance efficiency. Specifically, prior to initiating the BnB algorithm, we construct a regular grid, with a predefined width (set as $150$ in this study), over the initial cube $[\underline{\mathbf{r}}_{0},\overline{\mathbf{r}}_{0}]$. Subsequently, we compute the corresponding $R_{ij}$ values for all grid points and store them. During the execution of the BnB algorithm, when presented with a range $[\underline{\mathbf{r}},\overline{\mathbf{r}}]$, we first identify the grid points falling within this range. Subsequently, we determine the minimum and maximum precomputed $R_{ij}$ values for these points, effectively approximating $\min\{R_{ij}|\underline{\mathbf{r}}\leq\mathbf{r}\leq\overline{\mathbf{r}}\}$ and $\max\{R_{ij}|\underline{\mathbf{r}}\leq\mathbf{r}\leq\overline{\mathbf{r}}\}$.
8 Experiments
We implement the proposed algorithm, referred to as RPM-BP, using Matlab R2023b. To evaluate its performance, we conduct comparisons with other methods on a computer equipped with 6-core CPUs running at 3.2 GHz.
For the competing methods that solely produce point correspondences, we utilize the generated point correspondences to compute the optimal affine transformation between two sets of points. The matching error is then quantified as the root-mean-square distance between the transformed model inliers and their corresponding scene inliers.
To ensure efficient computation, we set the maximum branching depth of RPM-BP to 12.
8.1 2D registration
We conduct comparisons with RPM-HTB [6], RPM-PA [5] and RPM-CAV [4], all of which utilize global optimization techniques. These methods are adept at handling partial overlap and permit arbitrary similarity transformations between two point sets, rendering them suitable candidates for our comparative analysis.
8.1.1 2D synthetic data
Synthetic data offer a convenient means to assess specific facets of an algorithm’s performance. We conduct five distinct types of tests to gauge a method’s resilience to various disturbances: i) Deformation test: The prototype shape undergoes non-rigid deformation to generate the scene set. ii) Noise test: Positional noise is introduced to perturb the prototype shape, resulting in the scene point set. iii) Mixed outliers and inliers test: Random outliers are superimposed on the prototype shape to create both point sets. iv) Separate outliers and inliers test: Random outliers are added to different areas of the prototype shape, generating the two point sets. v) Occlusion+outlier test: Initially, the prototype shape is occluded to produce the two point sets. Subsequently, random outliers (with a fixed outlier-to-data ratio of 0.5) are incorporated into different regions of the point sets.
These tests are visually illustrated in Fig. 7. Additionally, to assess a method’s capability in handling arbitrary rotation and scaling, random rotations and scalings within the range of $[0.5,1.5]$ are applied when generating the two point sets. Fig. 7 provides examples of registrations performed by different methods.
The registration errors produced by various methods are depicted in the top two rows of Fig. 7. The result suggests that, in comparison to alternative methods, RPM-BP and RPM-HTB, with $n_{p}$ chosen as the ground truth, exhibits robustness against deformation, noise, and outliers when the outliers are distinct from the inliers. However, their robustness diminishes when outliers are intermixed with inliers. This limitation arises due to the fact that our method lacks $\epsilon-$global optimality, as discussed in Sec. 5.7, thereby affecting its robustness.
Regarding different transformation options, RPM-BP, when employing similarity transformation, demonstrates greater resilience to various disturbances than when utilizing affine transformation.
Furthermore, the choice of $n_{p}$ value significantly influences the performance of RPM-BP. Specifically, RPM-BP with an $n_{p}$ value close to the ground truth yields superior results. Moreover, RPM-BP displays less sensitivity to variations in the $n_{p}$ value compared to RPM-HTB, particularly evident in tests involving outliers and occlusion.
The average run times for various methods are depicted in the bottom row of Fig. 7. RPM-BP exhibits the best performance in scenarios involving deformation, positional noise, and mixed inlier-outlier tests. Comparing different transformation options, RPM-BP employing similarity transformation proves to be more efficient than when employing affine transformation. This efficiency is attributed to the lesser number of parameters involved in similarity transformation.
The outlier tests also provide insights into how various methods scale with problem size. Among them, RPM-BP, RPM-HTB, and RPM-CAV demonstrate the best scalability, followed by RPM-PA.
Refer to caption Refer to caption Refer to caption
Figure 5: a) to (c): Model point sets and examples of scene point sets in the deformation and noise tests, respectively. (d) to (i): Examples of model and scene point sets in the mixed outliers and inliers test ((d), (e)), separate outliers and inliers test ((f), (g)), and occlusion+outlier test ((h), (i)), respectively. In all cases, model points are indicated by red circles, while scene points are represented by blue crosses.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: Average registration errors (top 2 rows) and run times (bottom row) by RPM-BP, RPM-HTB [6], and RPM-CAV [4] under various $n_{p}$ values (ranging from $1/2$ to $/1$ 1 of the ground truth value) over 100 random trials for 2D deformation, noise, mixed outliers and inliers, separate outliers and inliers, and occlusion+outlier tests.
(a) RPM-BP (simi) (b) RPM-BP (aff) (c) RPM-HTB (simi) (d) RPM-HTB (aff) (e) RPM-PA (f) RPM-CAV
Figure 7: Example of registration results from different methods in the separate outliers and inliers test, with $n_{p}$ chosen as ground truth for all methods.
8.2 3D registration
8.2.1 3D synthetic data
Similar to the experimental setup detailed in Sec. 8.1.1, we conducted five types of tests to assess a method’s resilience against various disturbances. These tests include: i) Deformation test, ii) Noise test, iii) Mixed outliers and inliers test, iv) Separate outliers and inliers test, and v) Occlusion+outlier test. Refer to Fig. 8 for an illustration. Additionally, examples of registration by different methods are depicted in Fig. 10.
The top 2 rows of Fig. 9 displays registration errors from various methods. Results show that RPM-BP and RPM-HTB, with $n_{p}$ chosen as the ground truth, exhibits robustness against deformation, noise, and outliers when outliers are separate from inliers. However, they loses robustness when outliers are mixed with inliers. In contrast, GORE and TEASER++ struggle with handling deformation, noise, and outliers, especially when outliers are mixed with inliers, likely due to their reliance on features. In terms of different choices of $n_{p}$, RPM-BP performs significantly better when $n_{p}$ closely matches the ground truth.
The bottom row of Fig. 9 displays the average run times for various methods. TEASER++ emerges as the fastest, with Go-ICP following closely behind, and RPM-BP and RPM-HTB trailing after. GORE consistently exhibits slower performance across most cases. Notably, FRS experiences a rapid decline in efficiency when faced with challenging registration problems, such as an increase in outliers in both types of outlier tests.
Examining the two types of outlier tests also reveals how different methods scale with problem size. Go-ICP and GORE demonstrate the best scalability with problem size, followed by TEASER++, RPM-BP and RPM-HTB. Conversely, FRS exhibits the poorest scalability with problem size.
Refer to caption Refer to caption
Figure 8: (a) to (c): Model point sets and corresponding examples of scene point sets in the deformation and noise tests, respectively. (d) to (i): Examples of model and scene point sets in the mixed outliers and inliers test ((d), (e)), separate outliers and inliers test ((f), (g)), and occlusion+outlier test ((h), (i)), respectively. In all cases, model points are indicated by red circles, while scene points are represented by blue crosses.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Average registration errors (top 2 rows) and run times (bottom row) of RPM-BP, RPM-HTB, Go-ICP, FRS, GORE, and TEASER++ under varying $n_{p}$ values (ranging from $1/2$ to $/1$ 1 of the ground truth value) across 100 random trials for 3D deformation, noise, mixed outliers and inliers, separate outliers and inliers, and occlusion+outlier tests. For the separate outliers and inliers and occlusion+outlier tests, only a portion of FRS’s run time results is presented, as FRS experiences significant slowdowns with large problem sizes.
        (a) RPM-BP (b) RPM-HTB (c) Go-ICP (d) FRS (e) GORE (f) TEASER++
Figure 10: Examples of registration results from different methods in the separate outliers and inliers test, where the $n_{p}$ values of RPM-HTB and Go-ICP are both chosen to match the ground truth.
8.2.2 3DMatch dataset
The 3DMatch benchmark [41] encompasses scans of 62 scenes sourced from five established RGB-D reconstruction datasets (sun3d, 7-scenes, rgbd-scenes-v2, bundlefusion, and analysis-by-synthesis). D3Feat [42] offers a unified file structure and format. For convenience, we utilize the scan pairs provided by D3Feat to assess the performance of various methods.
The matching errors from different methods are depicted in Fig. 12. The results indicate that RPM-BP and RPM-HTB generally outperform Go-ICP and FRS but lag behind GORE and TEASER++. The superior performance of GORE and TEASER++ is partly attributed to their use of features, which significantly enhances their capability to match complex structures. In contrast, other methods rely solely on point position information. Additionally, the 3DMatch benchmark comprises purely rigid transformations, aligning well with the assumptions of GORE and TEASER++. Among RPM-BP and RPM-HTB, RPM-BP exhibits slightly superior performance, particularly evident in the sun3d and rgbd-scenes-v2 tests.
Examples of registration by different methods are presented in Fig. 12.
Refer to caption Refer to caption
Figure 11: Average registration errors from various methods across five RGB-D reconstruction datasets.
(a) RPM-BP (b) RPM-HTB (c) Go-ICP (d) FRS (e) GORE (f) TEASER++
Figure 12: Examples of registration results generated by different methods on the five RGB-D reconstruction datasets, arranged from top to bottom: sun3d, 7-scenes, rgbd-scenes-v2 (repeated), and analysis-by-synthesis.
9 Conclusion
In this paper, we introduced a BnB-based point set registration algorithm capable of aligning partially overlapping point sets while remaining invariant to the corresponding transformation. The method offers several advantages: the lower bound can be efficiently computed via linear assignment and low-dimensional convex quadratic program, and the dimensionality of the branching space matches the number of transformation parameters. These attributes make the method computationally efficient and scalable with problem size. By employing the BnB algorithm, our method demonstrates robustness against non-rigid deformations, positional noise, and outliers, particularly when outliers are distinct from inliers compared to state-of-the-art approaches.
However, the proposed method is limited to transformations with few parameters due to the exponential complexity of the BnB algorithm with respect to the dimensionality of the branching space. Additionally, the method’s robustness is significantly affected by the parameter setting for $n_{p}$. In the future, we plan to develop an adaptive scheme for setting $n_{p}$.
ACKNOWLEDGMENTS
This work was supported by the Fundamental Research Program of Shanxi Province (202103021223381, 202303021222267 and 202303021222271) and Science and Technological Innovation Program of Higher Education Institutions in Shanxi under Grant (2023L335)
Appendix A A weak result about positive semidefiniteness of matrix $\mathbf{H}^{0}+\mathbf{C}$
Proposition 1.
Matrix $\mathbf{H}^{0}+\mathbf{C}$ is close to being positive semidefinite.
Proof.
For any $\mathbf{p}\in\Omega$, it is apparently that $\mathbf{p}\geq 0$. Thus, given $\mathbf{p}\in\Omega$, the function $E$ in equation (3a) is a convex quadratic function of $\boldsymbol{\theta}$. Therefore, from equation (5), we can infer that $\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})+\mathbf{C}\succeq 0$ for any $\mathbf{p}\in\Omega$.
Let
$\mathbf{p}^{ij}=\arg\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$
Apparently ${\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p}^{ij})}+\mathbf{C}\succeq 0$. Based on the fact that $\underline{\Gamma}_{ij}=\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}=\min_{\mathbf{p}\in\{\mathbf{p}^{ij}\}}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$, we have that $\underline{\boldsymbol{\Gamma}}$ is close to one of $\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p}^{ij})$, $\forall i,j$. Consequently $\underline{\boldsymbol{\Gamma}}+\mathbf{C}$ is close to being positive semidefinite.
In an analogous way, let
$\mathbf{q}^{ij}=\arg\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$
Apparently ${\text{mat}(\mathbf{K}\mathbf{B}_{2}q^{ij})}+\mathbf{C}\succeq 0$. Based on $\overline{\Gamma}_{ij}=\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}=\max_{\mathbf{p}\in\{\mathbf{q}^{ij}\}}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$, we have $\underline{\boldsymbol{\Gamma}}$ is close to one of $\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{q}^{ij})$, $\forall i,j$. Consequently $\overline{\boldsymbol{\Gamma}}+\mathbf{C}$ is close to being positive semidefinite.
Therefore $\mathbf{H}^{0}+\mathbf{C}=\frac{\left(\underline{\boldsymbol{\Gamma}}+\mathbf{C}\right)+\left(\overline{\boldsymbol{\Gamma}}+\mathbf{C}\right)}{2}$ is close to being positive semidefinite.

### v2

@@
-Variable Substitution and Bilinear Programming for Aligning Partially Overlapping Point Sets
+Decomposed Global Optimization for Robust Point Matching with Low-Dimensional Branching
 Abstract
-In many applications, the demand arises for algorithms capable of aligning partially overlapping point sets while remaining invariant to the corresponding transformations.
-This research presents a method designed to meet such requirements through minimization of the objective function of the robust point matching (RPM) algorithm.
-First, we show that the RPM objective is a cubic polynomial.
-Then, through variable substitution, we transform the RPM objective to a quadratic function.
-Leveraging the convex envelope of bilinear monomials, we proceed to relax the resulting objective function, thus obtaining a lower bound problem that can be conveniently decomposed into distinct linear assignment and low-dimensional convex quadratic program components, both amenable to efficient optimization.
-Furthermore, a branch-and-bound (BnB) algorithm is devised, which solely branches over the transformation parameters, thereby boosting convergence rate.
-Empirical evaluations demonstrate better robustness of the proposed methodology against non-rigid deformation, positional noise, and outliers, particularly in scenarios where outliers remain distinct from inliers, when compared with prevailing state-of-the-art approaches.
+Numerous applications require algorithms that can align partially overlapping point sets while maintaining invariance to geometric transformations (e.g., similarity, affine, rigid).
+This paper introduces a novel global optimization method for this task by minimizing the objective function of the Robust Point Matching (RPM) algorithm.
+We first reveal that the original RPM objective is a cubic polynomial.
+Through a concise variable substitution, we transform this objective into a quadratic function.
+By leveraging the convex envelope of bilinear monomials, we derive a tight lower bound for this quadratic function.
+This lower bound problem conveniently and efficiently decomposes into two parts: a standard linear assignment problem (solvable in polynomial time) and a low-dimensional convex quadratic program.
+Furthermore, we devise a specialized Branch-and-Bound (BnB) algorithm that branches exclusively on the transformation parameters, which significantly accelerates convergence by confining the search space.
+Experiments on 2D and 3D synthetic and real-world data demonstrate that our method, compared to state-of-the-art approaches, exhibits superior robustness to non-rigid deformations, positional noise, and outliers, particularly in scenarios where outliers are distinct from inliers.
 keywords:
 branch-and-bound, partial overlap , bilinear monomial , point set registration, convex envelope, linear assignment
-journal: Neurocomputing
 1 Introduction
-Aligning 2D/3D point sets is crucial across computer vision, robotics, and remote sensing applications.
-Apart from disturbances like non-rigid deformation and positional noise, this task becomes especially challenging when the two point sets only partial overlap and lack initial coarse alignment.
-A practical approach to addressing the point set registration challenge involves minimizing the RPM objective [1].
-This objective is characterized by a mixed linear assignment-least squares formulation:
-$E(P,\boldsymbol{\theta})=\sum_{i,j}p_{ij}\|\mathbf{y}_{j}-\mathbf{T}(\mathbf{x}_{i}|\boldsymbol{\theta})\|^{2}$
-Here, ${\mathbf{x}_{i}}$ and ${\mathbf{y}_{j}}$ represent two point sets intended for alignment, $p_{ij}\in{0,1}$ indicates point correspondence, and $\mathbf{T}$ denotes a spatial transformation parameterized by $\boldsymbol{\theta}$.
-Lian et al.
-demonstrated that in scenarios where one point set can be embedded within another, after spatial transformations are eliminated, the RPM objective can be reduced to a concave quadratic function of point correspondence, exhibiting a low-rank structure [2].
-This simplification facilitates the implementation of efficient branch-and-bound (BnB) based global optimization strategies.
-The algorithm introduced in [2] is extended to handle partial overlap in [3].
-Here, the objective function takes the form of a non-quadratic concave function of point correspondence variable, exhibiting a low-rank structure.
-This structure enables application of efficient BnB optimization techniques.
-However, the algorithm’s branching space is a projected space of point correspondence, characterized by a dimensionality surpassing that of the transformation parameters.
-This results in a slower convergence rate for the method.
-Additionally, to ensure concavity of the objective function over the initial bounding region, the method requires the regularization of transformations, which in turn demands prior knowledge of transformation parameters.
-Consequently, the method lacks invariance to corresponding transformations, presenting a notable drawback.
+Point cloud registration, the process of aligning two point sets into a unified coordinate system, is a cornerstone of 3D computer vision, robotics, and augmented reality.
+While traditional methods like the Iterative Closest Point (ICP) algorithm are widely used, they’re notoriously sensitive to initial alignment and frequently get trapped in local minima, especially when dealing with significant outliers or partial data overlap.
+This fundamental limitation highlights the need for globally optimal algorithms that can reliably converge to the correct solution regardless of initialization.
+The Branch-and-Bound (BnB) framework has emerged as a powerful tool for achieving such global guarantees.
+Many work applies BnB to solve the coupled problem of finding the optimal correspondence $\mathbf{P}$ and transformation $\mathbf{T}(\cdot|\boldsymbol{\theta})$, typically formulated as minimizing the following objective function:
+$E(\mathbf{P},\boldsymbol{\theta})=\sum_{i,j}p_{ij}\|\mathbf{y}_{j}-\mathbf{T}(\mathbf{x}_{i}|\boldsymbol{\theta})\|^{2},$ (1)
+where $\mathbf{x}_{i}$} and $\mathbf{y}_{j}$} are points from the two sets, $p_{ij}\in\{0,1\}$ are binary variables indicating correspondence, and $\boldsymbol{\theta}$ parameterizes the spatial transformation.
+An Early attempt to solve this problem by branching over either $\mathbf{P}$ or $\boldsymbol{\theta}$ proved inefficient due to the immense search space and an inefficient bounding scheme.
+Subsequent research has largely split into two camps:
+  1. 1.
+    Correspondence-only Methods: These approaches eliminate the transformation variables $\boldsymbol{\theta}$, simplifying Eq.
+(1) into a problem that only involves correspondences.
+While this can reveal low-rank concave structures amenable to BnB optimization [1, 2], it comes at a cost.
+Initial methods were either restricted to subset alignment scenarios [1] or lacked transformation invariance [2].
+While follow-up works [3, 4] successfully restored invariance, they could not overcome the core limitation of this approach: branching in the high-dimensional correspondence space.
+This large search space inherently leads to slow convergence, making these methods impractical for larger problems.
+  2. 2.
+    Transformation-space Methods: This more promising strategy retains the transformation parameters $\boldsymbol{\theta}$ and performs branching within the much smaller, low-dimensional transformation space.
+A representative example is the work by [5], which treats Eq.
+(1) as a cubic polynomial and uses trilinear relaxation to derive a lower bound for the BnB algorithm.
+This approach effectively handles partial overlap and maintains transformation invariance, benefiting from a significantly reduced search space.
+1.1 Our Contributions
+In this paper, we follow the second strategy of branching on the low-dimensional transformation space.
+However, we introduce a novel method to derive a simpler, yet more effective, lower bound that avoids a key complexity of prior work.
+The algorithm in [5] requires a complex reformulation of the objective function to ensure one of the variables in the trilinear monomial includes the origin within its range.
+In contrast, as illustrated in Fig. 1, our method exclusively uses bilinear relaxation, which eliminates this requirement and results in a more straightforward and streamlined algorithm.
+Our main contributions are summarized as follows:
+  • We present a novel BnB algorithm for point cloud registration that utilizes a simpler, more robust lower-bounding scheme based on bilinear relaxation.
+This eliminates the need for complex objective function reformulations required by previous method [5], simplifying the overall algorithm.
+  • Our lower bound optimization is efficient.
+It decomposes into two well-behaved subproblems: a classic linear assignment problem for correspondences and a low-dimensional convex quadratic program for the transformation.
+Both can be solved to global optimality very quickly.
+  • Our resulting BnB algorithm inherits the key advantages of its predecessors, including robustness to partial overlap and invariance to transformations.
+Crucially, it leverages a low-dimensional branching space, enabling it to scale effectively to large problems while maintaining global optimality guarantees.
 […]
 Figure 1: Derivation of a lower bound function of the RPM objective function via variable substitution and bilinear relaxation.
-In response to this challenge, Lian and Zhang utilized the inherent constraints of rigid transformations to devise a novel objective function for the point correspondence variable.
-This function exhibits concavity across the entire solution space, eliminating the need for regularization of transformations [4].
-Alternatively, Lian et al.
-employed the polyhedral annexation (PA) algorithm to minimize the RPM objective [5].
-PA offers an advantage over BnB by operating exclusively within the concavity region of the objective, thus eliminating the need for transformation regularization.
-However, both approaches [4] and [5] still face the challenge inherited from [3] of branching within a high-dimensional projected space of point correspondence, resulting in sluggish convergence rates.
-The methods mentioned above all share a common practice: they remove spatial transformations from the RPM objective, resulting in a function based solely on point correspondence variable.
-However, this simplification leads to a complex function, posing challenges such as high-dimensional branching spaces and intricate optimization strategies.
-An alternative approach, as demonstrated by [6], suggests retaining spatial transformations.
-In this method, trilinear and bilinear relaxation are employed to derive a lower bound function.
-This approach offers several advantages, including its ability to handle partial overlap, its invariance to corresponding transformations, and its utilization of a low-dimensional branching space.
-Following the approach outlined in [6], we introduce a new Branch-and-Bound (BnB) algorithm tailored for optimizing the RPM objective function. This algorithm boasts the following features:
-  • The algorithm proposed by [6] is complex because it necessitates reformulating the RPM objective function to ensure that one of the variables within the trilinear monomial includes the origin within its range.
-In contrast, as illustrated in Fig.
-1, our method exclusively utilizes bilinear relaxation to establish the lower bound function, eliminating the necessity to fulfill this condition.
-This results in a streamlined algorithm compared to [6].
-  • The optimization of the lower bound function can be broken down into two distinct tasks: linear assignment and low-dimensional convex quadratic optimization related to spatial transformation variables.
-Both of these tasks can be efficiently solved with global optimality assurances.
-  • Our approach inherits the advantageous characteristics outlined in [6], including its ability in handling partial overlap, invariance to corresponding transformations, and utilization of a low-dimensional branching space.
-The paper is structured as follows: We begin with a review of related work (Sec.
-2).
-Following that, we introduce the theoretical framework necessary for developing our lower bound functions (Sec.
-3).
-Next, we delve into our objective function (Sec.
-4 and optimization strategy (5).
-Subsequently, we apply the proposed algorithm to 2D and 3D registration (Sec.
-6 and Sec.
-7).
-Finally, we present our experimental results (Sec.
-8) and conclude the paper (Sec.
-9).
+The rest of the paper is organized as follows.
+Section 2 provides a review of related work.
+Section 3 introduces the theoretical framework for our lower bound.
+Section 4 details our objective function, and Section 5 describes the optimization strategy.
+We apply the algorithm to 2D and 3D registration in Sections 6 and 7, respectively.
+Finally, we present experimental results in Section 8 and conclude in Section 9.
 2 Related Work
-A comprehensive survey of point set registration methods is available in [7]. Here, we focus solely on works closely related to ours.
+A comprehensive survey of point set registration is available in [6]. Here, we focus on works most closely related to our own, distinguishing between local and global optimization strategies.
 2.1 Local Methods
+These methods are efficient but do not guarantee global optimality, making them susceptible to local minima.
 Simultaneous Pose and Correspondence.
-The widely known ICP method [8] iterates between recovering point correspondence and updating spatial transformation.
-Despite its efficiency, the discrete nature of point correspondence often traps the method in local minima.
-Zhang et al.
-enhanced the robustness of ICP by constraining the rotation angle at each iterative step [9].
-Liang introduced RGB information to the registration process of ICP [10].
-The RPM method [1] relaxes point correspondence to be fuzzily valued and employs deterministic annealing (DA) to gradually recover point correspondence.
-Instead of DA, Sofka et al.
-utilized the covariance matrix of the transformation parameters to determine the fuzziness of point correspondence, resulting in better robustness to missing or extraneous structures [11].
-Ma et al.
-employed L2E to robustly estimate transformation for non-rigid registration [12].
-Our method is related to the aforementioned methods by also modeling spatial transformation and correspondences.
-However, our objective function is globally minimized, which enhances our method’s robustness to disturbances.
+The venerable Iterative Closest Point (ICP) algorithm [7, 8] alternates between finding point correspondences and updating the spatial transformation.
+While efficient, its reliance on discrete assignments often leads it to get trapped in local minima.
+To enhance robustness, variations like that by Zhang et al.
+[9] constrain the rotation angle, while others incorporate additional information, such as RGB data [10].
+The Robust Point Matching (RPM) method [11] addresses this by relaxing correspondences to have fuzzy values, using deterministic annealing (DA) to gradually refine them.
+Sofka et al.
+[12] used the covariance of transformation parameters to control this fuzziness, achieving better robustness to missing structures.
+For non-rigid registration, Ma et al.
+[13] employed L2E for robust transformation estimation.
+Our method, like these approaches, jointly models spatial transformation and correspondence.
+However, we achieve greater robustness to various disturbances by guaranteeing global optimality, which is not possible with these local methods.
 Correspondence-free Methods.
-The CPD method [13] frames point matching as fitting a Gaussian mixture model (GMM) representing one point set to another.
-GMMREG [14] utilizes two GMMs to represent two point sets and minimizes the $L_{2}$-norm distance between them.
-The SVGMs method represents point sets using sparse Gaussian components [15].
-Efficiency in GMM-based methods is enhanced by employing filtering to solve the correspondence problem [16].
-The density variation problem of point sets is addressed in [17] by modeling scene structure.
-Hierarchical Gaussian mixture representation improves speed and accuracy of registration [18].
+These methods, such as Coherent Point Drift (CPD) [14], avoid explicit correspondence matching.
+CPD frames registration as fitting a Gaussian Mixture Model (GMM) from one point set to another.
+GMMREG [15] represents both point sets with GMMs and minimizes the L2 distance between them.
+Similarly, SVGMs [16] use sparse Gaussian components.
+GMM-based methods have been made more efficient with filtering techniques [17], while issues with point set density have been addressed by modeling scene structure [18] and using hierarchical representations [19].
+These methods are computationally intensive and can still suffer from local minima, whereas our approach offers a global optimality guarantee.
 2.2 Global Methods
-BnB based methods.
-The Branch-and-Bound (BnB) algorithm stands as a widely-used global optimization technique.
-Rooted in Lipschitz theory, BnB is applied to align 3D shapes [19], albeit with limitations regarding outliers or occlusions.
-By capitalizing on the structure of 3D rigid motions geometry, Go-ICP [20] globally optimizes the objective of ICP.
-GOGMA [21] achieves point set registration by aligning Gaussian Mixture Models (GMMs) constructed from the original point sets.
-Employing stereographic projection, the fast rotation search (FRS) method efficiently recovers 3D rotation between two point sets [22].
-Bayesian nonparametric mixture and a novel approach of tessellating rotation space are introduced in [23], leading to an efficient alignment algorithm.
-Our BnB-based method distinguishes itself from the aforementioned methods by also incorporating point correspondences.
-This feature empowers our method to better handle non-rigid deformation.
+These methods offer guarantees of finding the globally optimal solution, albeit at a higher computational cost.
+BnB-based Methods.
+The Branch-and-Bound (BnB) algorithm is a powerful global optimization technique.
+An early application based on Lipschitz theory [20] was sensitive to outliers.
+More recent methods leverage the specific geometry of transformations.
+Go-ICP [21] globally optimizes the ICP objective by exploring the structure of 3D rigid motions.
+GOGMA [22] registers point sets by aligning their corresponding GMMs.
+Other works, such as the Fast Rotation Search (FRS) [23], efficiently find 3D rotations, while a Bayesian approach [24] uses a novel tessellation of the rotation space.
+While these methods offer global guarantees, they often focus on specific problems (e.g., rigid motion) and do not handle correspondence explicitly, which can limit their applicability to non-rigid deformations.
+Our method differs by jointly and globally optimizing both transformation and correspondence, making it more versatile.
 Mismatch Removal Methods.
-An alternative avenue of research focuses on recovering spatial transformations given putative point correspondences.
-The fast global registration (FGR) method [24] optimizes a robust objective grounded in the duality between line processes and robust estimation.
-GORE [25] eliminates a large portion of outliers using geometric operations before invoking RANSAC.
-TEASER++ [26] employs a graph-theoretic approach to decouple rotation, scale, and translation estimation.
-It adopts a truncated least squares formulation for each subproblem, enabling effective mismatch removal.
-2.3 Deep learning methods
-Correspondence-free methods.
-The Lucas–Kanade algorithm aligns global features of point sets generated by a PointNet network in [27].
-This approach is refined in [28] by employing a decoder to convert the generated global features back into point sets.
-Ensuring fidelity, Chamfer distances between the decoded point sets and the original ones are minimized.
-Correspondence-based methods.
-The FCGF method [29] employs sparse high-dimensional convolutions to extract dense feature descriptors from point clouds.
-This method significantly outperforms many hand-crafted features and PointNet-based methods.
-Additionally, it surpasses a 3D convolution-based method [30].
-The deep global registration method [31] utilizes FCGF feature descriptors for point cloud registration.
-DCP [32] utilizes DGCNN [33] and Transformer [34] to extract features for each point set, with SVD employed to recover rigid transformation.
-PRNet [35] extends DCP to handle partially overlapping point sets iteratively.
-IDAM [36] considers both geometric and distance features during its point matching process.
-RIENet [37] calculates feature-to-feature correspondences with neighborhood consensus.
-Finally, [38] performs partial point set registration by generating matching credibility on the global semantic level and learning geometric features on the local structural level.
+An alternative research direction focuses on finding the correct spatial transformation from an initial set of putative correspondences that may contain many errors.
+The Fast Global Registration (FGR) method [25] optimizes a robust objective based on line processes.
+GORE [26] prunes outliers using geometric constraints before applying RANSAC.
+TEASER++ [27] uses a graph-theoretic approach to decouple the estimation of rotation, scale, and translation, formulating each as a truncated least squares problem.
+TR-DE [28] also decomposes the 6-DOF transformation to reduce computational complexity within a two-stage BnB search.
+TEAR [29] uses a truncated entry-wise absolute residual function for outlier robustness.
+These methods excel at handling high outlier rates but require an initial set of correspondences, which can be difficult to obtain robustly in highly cluttered or unstructured environments.
+Our method avoids this by simultaneously solving for both correspondence and transformation from scratch.
 3 Bilinear Convex Envelope Linearization
-The convex envelope of a bilinear monomial $xy$ subject to constraints $\underline{x}\leq x\leq\overline{x}$ and $\underline{y}\leq y\leq\overline{y}$ can be described as:
-$\begin{cases}(xy)_{l1}\triangleq\underline{x}y+x\underline{y}-\underline{x}\underline{y}\leq xy,\\ (xy)_{l2}\triangleq\overline{x}y+x\overline{y}-\overline{x}\overline{y}\leq xy\end{cases}$ (1)
-There are two ways of utilizing this formula when devising the lower bound problem: (1) Integration of the term $\max\{(xy)_{l1},(xy)_{l2}\}$ into the lower bound objective function.
-Nevertheless, this incorporation can lead to non-linearity of the resultant function.
-(2) Formulation of a subproblem in the format $\min\{\alpha|\alpha\geq(xy)_{l1},\alpha\geq(xy)_{l2}\}$ within the lower bound problem.
-However, this approach increases the complexity of constraints and raises concerns regarding the feasibility of the resulting problem for efficient optimization via combinatorial optimization algorithms.
-Hence, in alignment with the methodology proposed in [6], we advocate for using the average of the convex envelope components, as delineated in (2):
-$(xy)_{l}\triangleq\frac{1}{2}\sum_{i}(xy)_{li}\leq xy$ (2)
-as the lower bound function for $xy$. This choice has the advantage that the lower bound function is linear, thereby facilitating efficient optimization by combinatorial optimization algorithms.
-[6] also shows that for $(xy)_{l}$ to converge to $xy$, we only need to branch over one variable $x$ or $y$ that makes up the bilinear monomial $xy$.
-This is a crucial property and the good convergence of the proposed algorithm hings on this fact.
+Our approach to constructing a lower bound for the BnB algorithm relies on the convex envelope of a bilinear monomial $xy$.
+For variables $x$ and $y$ within a rectangular domain defined by $\underline{x}\leq x\leq\overline{x}$ and $\underline{y}\leq y\leq\overline{y}$, the convex envelope is the maximum of two linear functions:
+$(xy)_{l}=\max\left\{\underline{x}y+x\underline{y}-\underline{x}\underline{y},\quad\overline{x}y+x\overline{y}-\overline{x}\overline{y}\right\}\leq xy.$ (2)
+While this provides the tightest possible linear lower bound, using a ‘max‘ function within our objective would result in a non-linear problem, complicating the subproblem optimization.
+To maintain linearity and ensure computational efficiency, we use a simpler, linearized lower bound. We use the average of the two linear components from the convex envelope:
+$(xy)_{l}\triangleq\frac{1}{2}\left(\left(\underline{x}y+x\underline{y}-\underline{x}\underline{y}\right)+\left(\overline{x}y+x\overline{y}-\overline{x}\overline{y}\right)\right)\leq xy.$ (3)
+This choice results in a strictly linear function, which is crucial for decomposing our lower bound subproblem into solvable components.
+A key property of this average-based lower bound is its relationship to the branching process.
+As shown by [5], for the lower bound to converge to the true value of the monomial, it is sufficient to branch on only one of the variables—either $x$ or $y$.
+This property is fundamental to the efficiency of our algorithm, as it allows us to effectively reduce the search space by branching only on the transformation parameters $\boldsymbol{\theta}$ while leaving the correspondence variables $p_{ij}$ unbranched.
+This single-variable branching strategy is the cornerstone of our algorithm’s convergence and its ability to scale to larger problems.
 4 Problem formulation
-Suppose there are two point sets $\mathscr{X}=\{\mathbf{x}_{i},1\leq i\leq n_{x}\}$ and $\mathscr{Y}=\{\mathbf{y}_{j},1\leq j\leq n_{y}\}$ in $\mathbb{R}^{n_{d}}$ to be aligned, where the column vectors $\mathbf{x}_{i}$ and $\mathbf{y}_{j}$ represent the coordinates of points $i$ and $j$, repectively.
-Based on [3], this problem can be modeled as a mixed linear assignment--least square problem:
+We aim to find a globally optimal alignment between two point sets, $\mathscr{X}=\{\mathbf{x}_{i}\}_{i=1}^{n_{x}}$ and $\mathscr{Y}=\{\mathbf{y}_{j}\}_{j=1}^{n_{y}}$ in $\mathbb{R}^{n_{d}}$.
+Following the mixed linear assignment-least squares framework of [2], the problem is formulated as minimizing the following objective function:
 $\min E(\mathbf{P},\boldsymbol{\theta})=\sum_{i,j}p_{ij}\|\mathbf{y}_{j}-\mathbf{J}(\mathbf{x}_{i})\boldsymbol{\theta}\|^{2}$ (4a)
 $={\boldsymbol{\theta}}^{\top}\mathbf{J}^{\top}(\text{diag}(\mathbf{P}\mathbf{1}_{n_{y}})\otimes\mathbf{I}_{d})\mathbf{J}{\boldsymbol{\theta}}-2{\boldsymbol{\theta}}^{\top}\mathbf{J}^{\top}(\mathbf{P}\otimes\mathbf{I}_{d})\mathbf{y}+\mathbf{1}_{n_{x}}^{\top}\mathbf{P}\widetilde{\mathbf{y}}$ (4b)
-[…]
-where
-$\boldsymbol{\theta}$ : parameters of the spatial transformation $\mathbf{T}(\mathbf{x}_{i}|\boldsymbol{\theta})=\mathbf{J}(\mathbf{x}_{i})\boldsymbol{\theta}$.
-$\mathbf{J}(\mathbf{x}_{i})$ : the Jacobian matrix at point $\mathbf{x}_{i}$.
-$\mathbf{I}_{d}$: the $d\times d$ identity matrix.
-$\mathbf{1}_{n_{x}}$: the $n_{x}$-dimensional vector of all ones.
-$\otimes$: the Kronecker product.
-$\mathbf{P}=\{p_{ij}\}$ : assignment matrix, with element $p_{ij}{=}1$ or 00 indicating that $\mathbf{x}_{i}$ matches $\mathbf{y}_{j}$ or not.
-$\mathbf{1}_{n_{x}}^{\top}\mathbf{P}\mathbf{1}_{n_{y}}{=}n_{p}$: the number of matches is a priori known to be a constant integer $n_{p}$.
+$s.t.\ \mathbf{P}\mathbf{1}_{n_{y}}\leq\mathbf{1}_{n_{x}},\ \mathbf{1}_{n_{x}}^{\top}\mathbf{P}\leq\mathbf{1}_{n_{y}}^{\top},\ \mathbf{1}_{n_{x}}^{\top}\mathbf{P}\mathbf{1}_{n_{y}}=n_{p},\ \mathbf{P}\geq 0,\ \underline{\boldsymbol{\theta}}\leq\boldsymbol{\theta}\leq\overline{\boldsymbol{\theta}}$ (4c)
+where $\mathbf{P}=\{p_{ij}\}$ is the binary assignment matrix, with $p_{ij}=1$ if point $\mathbf{x}_{i}$ matches $\mathbf{y}_{j}$, and 0 otherwise.
+The spatial transformation is parameterized by $\boldsymbol{\theta}$ such that $\mathbf{T}(\mathbf{x}_{i}|\boldsymbol{\theta})=\mathbf{J}(\mathbf{x}_{i})\boldsymbol{\theta}$, where $\mathbf{J}(\mathbf{x}_{i})$ is the Jacobian matrix at point $\mathbf{x}_{i}$.
+The constraint $\mathbf{1}_{n_{x}}^{\top}\mathbf{P} \mathbf{1}_{n_{y}}=n_{p}$ specifies that the number of matches is a known constant $n_{p}$, which models partial overlap.
+$\mathbf{I}_{d}$ is the $d\times d$ identity matrix.
+$\mathbf{1}_{n_{x}}$ is the $n_{x}$-dimensional vector of all ones.
+$\otimes$ is the Kronecker product.
 $\underline{\boldsymbol{\theta}}$ (resp.
-$\overline{\boldsymbol{\theta}}$) : the lower (resp.
+$\overline{\boldsymbol{\theta}}$) is the lower (resp.
 upper) bound of $\boldsymbol{\theta}$.
-$\mathbf{y}\triangleq\begin{bmatrix}\mathbf{y}_{1}^{\top},\dots,\mathbf{y}_{n_{y}}^{\top}\end{bmatrix}^{\top}$: a vector.
-$\widetilde{\mathbf{y}}\triangleq\begin{bmatrix}\|\mathbf{y}_{1}\|_{2}^{2},\dots,\|\mathbf{y}_{n_{y}}\|_{2}^{2}\end{bmatrix}^{\top}$: a vector.
-$\mathbf{J}\triangleq\begin{bmatrix}\mathbf{J}^{\top}(\mathbf{x}_{1}),\dots,\mathbf{J}^{\top}(\mathbf{x}_{n_{x}})\end{bmatrix}^{\top}$: a matrix.
-For ease of derivation in the sequel, we need to convert matrix $\mathbf{P}$ into a vector $\mathbf{p}$. Accordingly, the objective function takes the following form [3]:
+Vectors $\mathbf{y}\triangleq\begin{bmatrix}\mathbf{y}_{1}^{\top},\dots,\mathbf{y}_{n_{y}}^{\top}\end{bmatrix}^{\top}$, $\widetilde{\mathbf{y}}\triangleq\begin{bmatrix}\|\mathbf{y}_{1}\|_{2}^{2},\dots,\|\mathbf{y}_{n_{y}}\|_{2}^{2}\end{bmatrix}^{\top}$ and matrix $\mathbf{J}\triangleq\begin{bmatrix}\mathbf{J}^{\top}(\mathbf{x}_{1}),\dots,\mathbf{J}^{\top}(\mathbf{x}_{n_{x}})\end{bmatrix}^{\top}$.
+For the purpose of optimization, we reformulate the objective by vectorizing the matrices.
+By converting the assignment matrix $\mathbf{P}$ to a vector $\mathbf{p}=\text{vec}(\mathbf{P})$, the objective function can be written in a more compact form:
 $E(\mathbf{p},\boldsymbol{\theta})=\boldsymbol{\theta}^{\top}\text{mat}(\mathbf{B}\mathbf{p})\boldsymbol{\theta}-2\boldsymbol{\theta}^{\top}\mathbf{A}\mathbf{p}+\boldsymbol{\rho}^{\top}\mathbf{p}$ (5)
-where
-$\mathbf{p}\triangleq\text{vec}(\mathbf{P})$: a vector.
-$\boldsymbol{\rho}\triangleq\mathbf{1}_{n_{x}}\otimes\widetilde{\mathbf{y}}$: a vector.
-$\mathbf{A}\triangleq(\mathbf{J}^{\top}\otimes\mathbf{y}^{\top})\mathbf{W}^{n_{x},n_{y}}_{n_{d}}$: a matrix.
-$\mathbf{B}\triangleq(\mathbf{J}_{2}^{\top}\otimes\mathbf{I}_{n_{\theta}})\mathbf{W}^{n_{x},1}_{n_{\theta}}(\mathbf{I}_{n_{x}}\otimes\mathbf{1}_{n_{y}}^{\top})$: a matrix.
-$\mathbf{J}_{2}\triangleq\begin{bmatrix}\mathbf{J}(\mathbf{x}_{1})^{\top}\mathbf{J}(\mathbf{x}_{1}),\ldots,\mathbf{J}(\mathbf{x}_{n_{x}})^{\top}\mathbf{J}(\mathbf{x}_{n_{x}})\end{bmatrix}^{\top}$: a matrix.
-$\mathbf{W}^{m,n}_{d}\triangleq\mathbf{I}_{m}\otimes\begin{bmatrix}\mathbf{I}_{n}\otimes(\mathbf{e}_{d}^{1})^{\top},\dots,\mathbf{I}_{n}\otimes(\mathbf{e}_{d}^{d})^{\top}\end{bmatrix}^{\top}$: a matrix.
-$\text{vec}(\cdot)$: vectorization of a matrix by concatenating rows.
-$n_{\theta}$: the dimensionality of ${\boldsymbol{\theta}}$.
-$\mathbf{e}_{d}^{i}$: the $d$-dimensional column vector with only one nonzero unit element at the $i$-th position.
-$\text{mat}(\cdot)$: reconstructs a matrix from a vector, which is the inverse of the operator $\text{vec}(\cdot)$.
+Here, $\mathbf{A}\triangleq(\mathbf{J}^{\top}\otimes\mathbf{y}^{\top})\mathbf{W}^{n_{x},n_{y}}_{n_{d}}$, $\mathbf{B}\triangleq(\mathbf{J}_{2}^{\top}\otimes\mathbf{I}_{n_{\theta}})\mathbf{W}^{n_{x},1}_{n_{\theta}}(\mathbf{I}_{n_{x}}\otimes\mathbf{1}_{n_{y}}^{\top})$, and $\boldsymbol{\rho}\triangleq\mathbf{1}_{n_{x}}\otimes\widetilde{\mathbf{y}}$ are constant matrices and vectors derived from the point coordinates and Jacobian matrices, Matrices $\mathbf{J}_{2}\triangleq\begin{bmatrix}\mathbf{J}(\mathbf{x}_{1})^{\top}\mathbf{J}(\mathbf{x}_{1}),\ldots,\mathbf{J}(\mathbf{x}_{n_{x}})^{\top}\mathbf{J}(\mathbf{x}_{n_{x}})\end{bmatrix}^{\top}$ and $\mathbf{W}^{m,n}_{d}\triangleq\mathbf{I}_{m}\otimes\begin{bmatrix}\mathbf{I}_{n}\otimes(\mathbf{e}_{d}^{1})^{\top},\dots,\mathbf{I}_{n}\otimes(\mathbf{e}_{d}^{d})^{\top}\end{bmatrix}^{\top}$.
+$n_{\theta}$ denotes the dimensionality of $\boldsymbol{\theta}$}.
+$\mathbf{e}_{d}^{i}$ is the $d$-dimensional column vector with only one nonzero unit element at the $i$-th position.
+$\text{mat}(\cdot)$ reconstructs a matrix from a vector, which is the inverse of the operator $\text{vec}(\cdot)$.
 Since $\mathbf{1}^{\top}_{n_{x}n_{y}}\mathbf{p}=n_{p}$, a constant, for rows of $\mathbf{B}$ containing identical elements, the result of them multiplied by $\mathbf{p}$ can be replaced by constants.
@@
 where the nonzero elements of the constant matrix $\mathbf{C}$ correspond to the rows of $\mathbf{B}$ containing identical elements.
+(Please refer to Sec.
+6 and 7 for examples of $\mathbf{C}$).
 The elements of the constant matrix $\mathbf{K}$ take values in $\{0,1\}$ and record the correspondences between the rows of $\mathbf{B}_{2}$ and those of $\mathbf{B}$.
@@
 5 Optimization
-In this section, we begin by applying the bilinear relaxation formula, introduced in Section 3, to derive a lower bound function for our objective.
-This lower bound function’s optimization can then be divided into separate optimizations over $\boldsymbol{\theta}$ and $\mathbf{p}$ (discussed in Section 5.1).
-Subsequently, we demonstrate that it suffices to branch solely over $\boldsymbol{\theta}$ (explained in Section 5.2) and introduce a more effective method for computing the upper bound (outlined in Section 5.4).
-Lastly, we introduce the BnB algorithm (presented in Section 5.5), which offers global optimization for the objective function.
+This section details the optimization strategy for our Branch-and-Bound (BnB) algorithm.
+We begin by deriving a lower bound for our objective function using bilinear relaxation (Sec.
+5.1).
+We show that the minimization of this lower bound can be efficiently decoupled into two independent subproblems.
+Next, we present our branching strategy (Sec.
+5.2), demonstrating that it’s sufficient to branch on only the low-dimensional transformation parameters.
+Finally, we discuss how to compute a more effective upper bound (Sec.
+5.4) and outline the complete BnB algorithm (Sec.
+5.5).
 5.1 Lower bound function derivation
@@
 $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$), making optimization challenging.
-To tackle this challenge, we use the bilinear relaxation formula introduced in Section 3 to relax these terms into linear ones.
+To tackle this challenge, we use the bilinear relaxation formula introduced in Sec.
+3 to relax these terms into linear ones.
 The first relaxation is:
 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boxed{\text{trace}(\boldsymbol{\Gamma}\boldsymbol{\Theta})}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boxed{\text{trace}(\mathbf{H}^{0}\boldsymbol{\Theta})+\text{trace}(\mathbf{H}^{1}\boldsymbol{\Gamma})+\text{trace}(\mathbf{H}^{2}\mathbf{1}\mathbf{1}^{\top})}$}}bilinearrelaxation (8)
-where in accordance with equations (1) and (2), the coefficients can be calculated as
+where in accordance with equation (3), the coefficients can be calculated as
 $\mathbf{H}^{0}=\frac{\underline{\boldsymbol{\Gamma}}+\overline{\boldsymbol{\Gamma}}}{2},\mathbf{H}^{1}=\frac{\underline{\boldsymbol{\Theta}}+\overline{\boldsymbol{\Theta}}}{2},\mathbf{H}^{2}=-\frac{\underline{\boldsymbol{\Gamma}}\underline{\boldsymbol{\Theta}}+\overline{\boldsymbol{\Gamma}}\overline{\boldsymbol{\Theta}}}{2}$ (9)
@@
 upper) bounds of matrices $\boldsymbol{\Theta}$ and $\boldsymbol{\Gamma}$.
-It is also worth noting that the range $[\underline{\boldsymbol{\Theta}}$, $\overline{\boldsymbol{\Theta}}]$ of $\boldsymbol{\Theta}$ can be derived from the range $[\underline{\boldsymbol{\theta}},\overline{\boldsymbol{\theta}}]$ of $\boldsymbol{\theta}$ based on relation $\boldsymbol{\Theta}=\boldsymbol{\theta}\boldsymbol{\theta}^{\top}$, as explained in Section 5.3.
+It is also worth noting that the range $\underline{\boldsymbol{\Theta}}$, $\overline{\boldsymbol{\Theta}}]$ of $\boldsymbol{\Theta}$ can be derived from the range $\underline{\boldsymbol{\theta}},\overline{\boldsymbol{\theta}}]$ of $\boldsymbol{\theta}$ based on relation $\boldsymbol{\Theta}=\boldsymbol{\theta}\boldsymbol{\theta}^{\top}$, as explained in Sec.
+5.3.
 The second relaxation is:
 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boxed{\boldsymbol{\theta}^{\top}\boldsymbol{\eta}}$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boxed{\begin{matrix}(\mathbf{g}^{0})^{\top}\boldsymbol{\theta}+(\mathbf{g}^{1})^{\top}\boldsymbol{\eta}+\mathbf{1}_{n_{\theta}}^{\top}\mathbf{g}^{2}\end{matrix}}$}}bilinearrelaxation (10)
-where in accordance with equations (1) and (2), the coefficients can be calculated as
+where in accordance with equation (3), the coefficients can be calculated as
 $\mathbf{g}^{0}=\frac{\underline{\boldsymbol{\eta}}+\overline{\boldsymbol{\eta}}}{2},\mathbf{g}^{1}=\frac{\underline{\boldsymbol{\theta}}+\overline{\boldsymbol{\theta}}}{2},\mathbf{g}^{2}=\frac{\underline{\boldsymbol{\eta}}\underline{\boldsymbol{\theta}}+\overline{\boldsymbol{\eta}}\overline{\boldsymbol{\theta}}}{2}$ (11)
@@
 $+\boldsymbol{\theta}^{\top}\mathbf{C}\boldsymbol{\theta}+\boldsymbol{\theta}^{\top}\mathbf{g}^{0}+(\mathbf{g}^{1})^{\top}\boldsymbol{\eta}+\mathbf{1}^{\top}\mathbf{g}^{2}{+}\boldsymbol{\rho}^{\top}\mathbf{p}$ (12)
-By substituting back $\boldsymbol{\theta}\boldsymbol{\theta}^{\top}\Leftarrow\boldsymbol{\Theta}$, $\text{mat}(K\mathbf{B}_{2}\mathbf{p})\Leftarrow\boldsymbol{\Gamma}$ and $-2\mathbf{A}\mathbf{p}\Leftarrow\boldsymbol{\eta}$, we get the final form of the lower bound function:
+By substituting back $\boldsymbol{\theta}\boldsymbol{\theta}^{\top}\Leftarrow\boldsymbol{\Theta}$, $\text{mat}(K\mathbf{B}_{2}\mathbf{p})\Leftarrow\boldsymbol{\Gamma}$ and $2\mathbf{A}\mathbf{p}\Leftarrow\boldsymbol{\eta}$, we get the final form of the lower bound function:
 $E_{l}(\mathbf{p},\boldsymbol{\theta})=$ $\boldsymbol{\theta}^{\top}\mathbf{H}^{0}\boldsymbol{\theta}+\text{trace}(\mathbf{H}^{1}\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p}))+\text{trace}(\mathbf{H}^{2}\mathbf{1}\mathbf{1}^{\top})$
@@
 It is apparent that the minimization of $E_{l}$ under constraints (4c) can be decomposed into separate optimizations over $\mathbf{p}$ and $\boldsymbol{\theta}$.
-The optimization over $\mathbf{p}$ is
+1. Optimization over $\mathbf{p}$. This subproblem is a linear program in terms of the correspondence vector $\mathbf{p}$.
 $\min_{p}\ \text{trace}(\mathbf{H}^{1}\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p}))+(\mathbf{g}^{1})^{\top}(-2\mathbf{A}\mathbf{p})+\boldsymbol{\rho}^{\top}\mathbf{p}$ (14a)
@@
 This is an $n_{p}$-cardinality linear assignment problem and can be transformed into a standard linear assignment problem [30] and then efficiently solved by combinatorial optimization algorithms such as LAPJV [31].
-The optimization over $\boldsymbol{\theta}$ is:
+2. Optimization over $\boldsymbol{\theta}$. This subproblem is a quadratic program in terms of the transformation vector $\boldsymbol{\theta}$.
 $\min_{\boldsymbol{\theta}}\ \boldsymbol{\theta}^{\top}(\mathbf{H}^{0}+\mathbf{C})\boldsymbol{\theta}+\boldsymbol{\theta}^{\top}\mathbf{g}^{0}$ (15a)
@@
 5.2 Branching strategy
-According to the result stated in Section 3, for the bilinear term $\boldsymbol{\theta}^{\top}\boldsymbol{\eta}$, for its linear lower bound function to converge to it, it suffices to branch over $\boldsymbol{\theta}$ and leave the range of $\boldsymbol{\eta}$ fixed.
-Here, utilizing $\boldsymbol{\eta}=-2\mathbf{A}\mathbf{p}$, the fixed range $[\underline{\boldsymbol{\eta}},\overline{\boldsymbol{\eta}}]$ of $\boldsymbol{\eta}$ can be computed as
-$\underline{\eta}_{i}=\min_{\mathbf{p}\in\Omega}(-2\mathbf{A}\mathbf{p})_{i}$ (15)
-$\overline{\eta}_{i}=\max_{\mathbf{p}\in\Omega}(-2\mathbf{A}\mathbf{p})_{i}$ (16)
+According to the result stated in Sec.
+3, for the bilinear term $\boldsymbol{\theta}^{\top}\boldsymbol{\eta}$, for its linear lower bound function to converge to it, it suffices to branch over $\boldsymbol{\theta}$ and leave the range of $\boldsymbol{\eta}$ fixed.
+Here, utilizing $\boldsymbol{\eta}=-2\mathbf{A}\mathbf{p}$, the fixed range $\underline{\boldsymbol{\eta}},\overline{\boldsymbol{\eta}}]$ of $\boldsymbol{\eta}$ can be computed as
+$\underline{\eta}_{i}=\min_{\mathbf{p}\in\Omega}(-2\mathbf{A}\mathbf{p})_{i},\quad\overline{\eta}_{i}=\max_{\mathbf{p}\in\Omega}(-2\mathbf{A}\mathbf{p})_{i}$ (16)
 Likewise, for the bilinear term $\text{trace}(\boldsymbol{\Gamma}\boldsymbol{\Theta})$, for its linear lower bound function to converge to it, we only need to branch over $\boldsymbol{\theta}$ (since the range of $\boldsymbol{\theta}$ determines the range of $\boldsymbol{\Theta}$ as explained in Sec.
 5.3) and leave the range of $\boldsymbol{\Gamma}$ fixed.
-Here, utilizing $\boldsymbol{\Gamma}=\text{mat}(K\mathbf{B}_{2}\mathbf{p})$, the fixed range $[\underline{\boldsymbol{\Gamma}},\overline{\boldsymbol{\Gamma}}]$ of $\boldsymbol{\Gamma}$ can be computed as follows:
-$\underline{\Gamma}_{ij}=\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$ (17)
-$\overline{\Gamma}_{ij}=\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$ (18)
-With thess choice of $\underline{\boldsymbol{\Gamma}}$ and $\overline{\boldsymbol{\Gamma}}$, it has been empirically demonstrated that
+Here, utilizing $\boldsymbol{\Gamma}=\text{mat}(K\mathbf{B}_{2}\mathbf{p})$, the fixed range $\underline{\boldsymbol{\Gamma}},\overline{\boldsymbol{\Gamma}}]$ of $\boldsymbol{\Gamma}$ can be computed as follows:
+$\underline{\Gamma}_{ij}=\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij},\quad\overline{\Gamma}_{ij}=\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$ (17)
+With these choice of $\underline{\boldsymbol{\Gamma}}$ and $\overline{\boldsymbol{\Gamma}}$, it has been empirically demonstrated that
 $\mathbf{H}^{0}+\mathbf{C}=\frac{\left(\underline{\boldsymbol{\Gamma}}+\mathbf{C}\right)+\left(\overline{\boldsymbol{\Gamma}}+\mathbf{C}\right)}{2}\succeq 0$ (18)
-Please refer to Section 5.6 for relevant experiments and Appendix for a weaker assertion.
+Please refer to Sec.
+5.6 for relevant experiments and A for a weaker assertion.
 This result implies that problem (15a) (15b) is now a convex quadratic program.
@@
 We will address the proof of (18) in our future research.
-We also note that (15), (16), (17) and (18) are $n_{p}-$cardinlaity linear assignment problems, which can be efficiently solved by the aforementioned algorithms.
+We also note that (16) and (17) are $n_{p}$-cardinlaity linear assignment problems, which can be efficiently solved by the aforementioned algorithms.
 In conclusion, in our BnB algorithm, we only need to branch over $\boldsymbol{\theta}$.
@@
 Given the relationship $\boldsymbol{\Theta}=\boldsymbol{\theta}\boldsymbol{\theta}^{\top}$, and the specified range $\left[\underline{\boldsymbol{\theta}},\overline{\boldsymbol{\theta}}\right]$ of $\boldsymbol{\theta}$, we can determine the range $\left[\underline{\boldsymbol{\Theta}},\overline{\boldsymbol{\Theta}}\right]$ of $\boldsymbol{\Theta}$ as follows:
-$[\underline{\Theta_{ij}},\overline{\Theta_{ij}}]=\begin{cases}[\underline{\theta_{i}},\overline{\theta_{i}}]\cdot[\underline{\theta_{j}},\overline{\theta_{j}}],&\text{if}\ i\neq j\\ [\underline{\theta_{i}},\overline{\theta_{i}}]^{2},&\text{if}\ i=j\end{cases}$
+$\underline{\Theta_{ij}},\overline{\Theta_{ij}}]=\begin{cases}[\underline{\theta_{i}},\overline{\theta_{i}}]\cdot[\underline{\theta_{j}},\overline{\theta_{j}}],&\text{if}\ i\neq j\\ [\underline{\theta_{i}},\overline{\theta_{i}}]^{2},&\text{if}\ i=j\end{cases}$
 Here, employing interval arithmetic, the product operator is defined as:
-$[\underline{\theta_{i}},\overline{\theta_{i}}]\cdot[\underline{\theta_{j}},\overline{\theta_{j}}]=[\min(\underline{\theta_{i}}\underline{\theta_{j}},\underline{\theta_{i}}\overline{\theta_{j}},\overline{\theta_{i}}\underline{\theta_{j}},\overline{\theta_{i}}\overline{\theta_{j}}),\max(\underline{\theta_{i}}\underline{\theta_{j}},\underline{\theta_{i}}\overline{\theta_{j}},\overline{\theta_{i}}\underline{\theta_{j}},\overline{\theta_{i}}\overline{\theta_{j}})]$
+$\underline{\theta_{i}},\overline{\theta_{i}}]\cdot[\underline{\theta_{j}},\overline{\theta_{j}}]=[\min(\underline{\theta_{i}}\underline{\theta_{j}},\underline{\theta_{i}}\overline{\theta_{j}},\overline{\theta_{i}}\underline{\theta_{j}},\overline{\theta_{i}}\overline{\theta_{j}}),\max(\underline{\theta_{i}}\underline{\theta_{j}},\underline{\theta_{i}}\overline{\theta_{j}},\overline{\theta_{i}}\underline{\theta_{j}},\overline{\theta_{i}}\overline{\theta_{j}})]$
 Additionally, the square operator is defined as:
-$[\underline{\theta_{i}},\overline{\theta_{i}}]^{2}=\begin{cases}[\min(\underline{\theta_{i}}^{2},\overline{\theta_{i}}^{2}),\max(\underline{\theta_{i}}^{2},\overline{\theta_{i}}^{2})],&\text{if}\ 0\notin[\underline{\theta_{i}},\overline{\theta_{i}}]\\ [0,\max(\underline{\theta_{i}}^{2},\overline{\theta_{i}}^{2})],&\text{if}\ 0\in[\underline{\theta_{i}},\overline{\theta_{i}}]\end{cases}$
+[…]
 5.4 Upper bound
@@
 5.5 Branch-and-Bound
-Based on the aforementioned preparations, we are now poised to apply the BnB algorithm to optimize $E$.
-To initiate the process, we utilize the initial range of $\boldsymbol{\theta}$ to form the initial hypercube $M$.
-Subsequently, during each iteration of the algorithm, we identify the hypercube that yields the lowest lower bound among all available hypercubes.
-This selected hypercube is then further subdivided to enhance the global lower bound of the problem.
-Simultaneously, the upper bound is refined by assessing $E(\mathbf{p})$ using solutions generated during the computation of the lower bounds.
-For a comprehensive overview, refer to Algorithm 1.
+With the preceding preparations in place, we now apply the Branch-and-Bound (BnB) algorithm to optimize the objective function $E$.
+We begin by forming an initial hypercube $M$ using the initial range of the parameters $\boldsymbol{\theta}$.
+In each iteration, the algorithm selects the hypercube with the lowest lower bound for subdivision.
+This process progressively tightens the global lower bound.
+Simultaneously, the upper bound is refined by evaluating $E(\mathbf{p})$ using solutions derived during the lower bound computation.
+For a complete overview of the process, see Algorithm 1.
 1
@@
 5.6 Positive Semidefiniteness of $\mathbf{H}^{0}+\mathbf{C}$
-To confirm the assertion that $\mathbf{H}^{0}+\mathbf{C}\succeq 0$, we utilize the separate outliers and inliers test detailed in Section 8.1.1, with an outlier to data ratio set at 0.3.
-We present histograms of the smallest eigenvalues of the matrix $\mathbf{H}^{0}+\mathbf{C}$ across 100 registration instances in Figure 3.
-The result consistently support the claim that $\mathbf{H}^{0}+\mathbf{C}\succeq 0$, underscoring the efficacy of the proposed algorithm.
+To verify the positive semidefiniteness of the matrix $\mathbf{H}^{0}+\mathbf{C}$ ($\mathbf{H}^{0}+\mathbf{C}\succeq 0$), we conducted an experiment using the separate outliers and inliers test, as detailed in Section 8.1.1.
+The outlier-to-data ratio was set at 0.3.
+Figure 3 displays histograms of the smallest eigenvalues of $\mathbf{H}^{0}+\mathbf{C}$ across 100 different registration instances.
+The results consistently show that the smallest eigenvalue is non-negative, providing strong evidence that $\mathbf{H}^{0}+\mathbf{C}$ is indeed positive semidefinite.
+This outcome highlights the reliability and effectiveness of the proposed algorithm.
 Refer to caption Refer to caption
 Figure 3: Histograms depicting the smallest eigenvalue of $\mathbf{H}^{0}+\mathbf{C}$ across 100 registration instances.
-5.7 Convergence of the BnB algorithm
-To assess the convergence of the proposed BnB algorithm, we conducted the separate outliers and inliers test detailed in Sec.
-8.1.1, with an outlier to data ratio set at $0.3$.
-The experimental results are depicted in Fig.
-4.
-Key observations include: 1) A narrower initial range of $\boldsymbol{\theta}$ results in a tighter duality gap, evidenced by improved upper and lower bounds.
-This suggests that leveraging additional information about the range of $\boldsymbol{\theta}$ could effectively reduce the duality gap in practical applications.
-2) Easier problem instances, such as the fish test compared to the character test, exhibit a tighter duality gap, thus leads to faster convergence.
-Properties 2) suggests that employing a fixed duality gap threshold as the stopping criterion for all types of problems may not be a good idea.
-It may happen that a threshold is set too small for challenging problems, leading to prolonged execution times.
-Conversely, it might be set too large for simpler problems, resulting in premature termination.
-Therefore, following [6], we adopt the maximum branching depth as our stopping criterion.
+5.7 BnB Algorithm Convergence
+We assessed the convergence of our BnB algorithm using the separate outliers and inliers test (Sec. 8.1.1) with an outlier-to-data ratio of $0.3$. The results in Figure 4 reveal two key points:
+  1. 1.
+    A smaller initial range for $\boldsymbol{\theta}$ yields a tighter duality gap, which improves both upper and lower bounds.
+This suggests that leveraging a tighter initial range can speed up convergence.
+  2. 2.
+    Simpler problem instances (e.g., the fish test vs. the character test) have a tighter duality gap, leading to faster convergence.
+Based on observation 2, a fixed duality gap threshold is an unsuitable stopping criterion for all problem types, as it can cause either excessive computation time or premature termination.
+Therefore, consistent with [5], we use the maximum branching depth as our stopping criterion.
+Refer to caption Refer to caption
+Refer to caption Refer to caption
 Figure 4: Upper (first row) and lower bounds (second row) generated in each iteration of RPM-BP.
-Our method is tested with the $n_{p}$ value chosen as the ground truth and with varying initial ranges of $\boldsymbol{\theta}$: $[\boldsymbol{\theta}_{gt}-\Delta,\boldsymbol{\theta}_{gt}+\Delta]$, where $\boldsymbol{\theta}_{gt}$ represents the ground truth $\boldsymbol{\theta}$ solution.
-Here, the margin $\Delta$ takes values of $0.5$, $1$, and $1.5$ respectively.
-6 Application 1: 2D similarity/affine registration
-Our algorithm can handle 2D similarity and affine transformations since both of them can be written as the form of $T(\mathbf{x}|\boldsymbol{\theta})=J(\mathbf{x})\boldsymbol{\theta}$.
-A 2D similarity transformation has the form
-$T(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&-\theta_{2}\\ \theta_{2}&\theta_{1}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2}\end{bmatrix}+\begin{bmatrix}\theta_{3}\\ \theta_{4}\end{bmatrix}=\begin{bmatrix}x^{1}&-x^{2}&1&0\\ x^{2}&x^{1}&0&1\end{bmatrix}\boldsymbol{\theta}$
-where $[\theta_{3},\theta_{4}]^{\top}$ is translation and $\theta_{1}=s\cos(\phi)$ and $\theta_{2}=s\sin(\phi)$ with $s$ being scale and $\phi$ being rotation angle.
-It can be derived that $\mathbf{B}_{2}=\mathbf{B}([1,3,4],:)$ and $\mathbf{C}=\text{diag}([0,0,n_{p},n_{p}])$.
-A 2D affine transformation has the form
-$T(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&\theta_{2}\\ \theta_{3}&\theta_{4}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2}\end{bmatrix}+\begin{bmatrix}\theta_{5}\\ \theta_{6}\end{bmatrix}=\begin{bmatrix}x^{1}&x^{2}&0&0&1&0\\ 0&0&x^{1}&x^{2}&0&1\end{bmatrix}\boldsymbol{\theta}$
-It can be derived that $\mathbf{B}_{2}=\mathbf{B}([1,2,5,8,11],:)$ and $\mathbf{C}=\text{diag}([0,0,0,0,n_{p},n_{p}])$.
+Our method is tested with the $n_{p}$ value chosen as the ground truth and with varying initial ranges of $\boldsymbol{\theta}$: $\boldsymbol{\theta}_{gt}-\Delta,\boldsymbol{\theta}_{gt}+\Delta]$, where $\boldsymbol{\theta}_{gt}$ represents the ground truth $\boldsymbol{\theta}$ solution.
+Here, the margin $\Delta$ takes values of $0.5$, 11, and $1.5$ respectively.
+6 Application 1: 2D Similarity and Affine Registration
+Our algorithm is designed to handle various transformations as long as they can be expressed in the form $\mathbf{T}(\mathbf{x}|\boldsymbol{\theta})=\mathbf{J}(\mathbf{x})\boldsymbol{\theta}$.
+Both 2D similarity and affine transformations fit this structure, allowing for a straightforward application of our framework.
+For a 2D similarity transformation, the transformation matrix and parameter vector are:
+$\mathbf{T}(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&-\theta_{2}\\ \theta_{2}&\theta_{1}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2} \end{bmatrix}+\begin{bmatrix}\theta_{3}\\ \theta_{4}\end{bmatrix}=\begin{bmatrix}x^{1}&-x^{2}&1&0\\ x^{2}&x^{1}&0&1\end{bmatrix}\boldsymbol{\theta}$  (20)
+Here, $\theta_{3},\theta_{4}]^{\top}$ represents the translation, while $\theta_{1}$ and $\theta_{2}$ are derived from the scale $s$ and rotation angle $\phi$ as $\theta_{1}=s\cos(\phi)$ and $\theta_{2}=s\sin(\phi)$.
+For this transformation, it can be derived that $\mathbf{B}_{2}$ corresponds to rows $[1,3,4]$ of $\mathbf{B}$, and $\mathbf{C}=\text{diag}([0,0,n_{p},n_{p}])$.
+A 2D affine transformation is a more general case:
+$\mathbf{T}(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&\theta_{2}\\ \theta_{3}&\theta_{4}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2} \end{bmatrix}+\begin{bmatrix}\theta_{5}\\ \theta_{6}\end{bmatrix}=\begin{bmatrix}x^{1}&x^{2}&0&0&1&0\\ 0&0&x^{1}&x^{2}&0&1\end{bmatrix}\boldsymbol{\theta}$  (21)
+For this transformation, $\mathbf{B}_{2}$ corresponds to rows $[1,2,5,8,11]$ of $\mathbf{B}$, and $\mathbf{C}=\text{diag}([0,0,0,0,n_{p},n_{p}])$.
 7 Application 2: 3D Rigid Registration
-Theoretically, our method can handle 3D affine transformation since it takes the form $T(\mathbf{x}|\boldsymbol{\theta})=J(\mathbf{x})\boldsymbol{\theta}$. A 2D affine transformation has the form
-$T(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&\theta_{2}&\theta_{3}\\ \theta_{4}&\theta_{5}&\theta_{6}\\ \theta_{7}&\theta_{8}&\theta_{9}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2}\\ x^{3}\end{bmatrix}+\begin{bmatrix}\theta_{10}\\ \theta_{11}\\ \theta_{12}\end{bmatrix}$
-=$\begin{bmatrix}x^{1}&x^{2}&x^{3}&0&0&0&0&0&0&1&0&0\\ 0&0&0&x^{1}&x^{2}&x^{3}&0&0&0&0&1&0\\ 0&0&0&0&0&0&x^{1}&x^{2}&x^{3}&0&0&1\end{bmatrix}\boldsymbol{\theta}$
-It can be derived that $\mathbf{B}_{2}=\mathbf{B}([1,2,3,10,14,15,22,27,34],:)$ and $\mathbf{C}=\text{diag}([\underbrace{0,\ldots,0}_{9\text{ zeros}},n_{p},n_{p},n_{p}])$.
-Nevertheless, the 3D affine transformation has many parameters, causing our method to converge slowly.
-Therefore, for the 3D case, we instead focus on rigid transformation instead.
+Our method is theoretically capable of handling 3D affine transformations, which has the form
 […]
+It can be derived that $\mathbf{B}_{2}$ corresponds to rows $[1,2,3,10,14,15,22,27,34]$ of $\mathbf{B}$, and $\mathbf{C}=\text{diag}([\underbrace{0,\ldots,0}_{9\text{ zeros}},n_{p},n_{p},n_{p}])$.
+Nevertheless, the large number of parameters (12 for a full affine transformation) results in a high-dimensional branching space, significantly slowing down convergence.
+Therefore, for 3D point cloud registration, we focus on the more constrained rigid transformation.
+A 3D rigid transformation involves a rotation and a translation.
+We represent the 3D rotation using the angle-axis representation, where a rotation is defined by a 3D vector $\mathbf{r}$.
+The corresponding $3\times 3$ rotation matrix $\mathbf{R}\in\mathbb{SO}_{3}$ can be obtained via the matrix exponential map [21]:
 $\mathbf{R}=\mathbf{I}_{3}+\frac{[\mathbf{r}]_{\times}\sin\|\mathbf{r}\|}{\|\mathbf{r}\|}+\frac{[\mathbf{r}]_{\times}^{2}(1-\cos\|\mathbf{r}\|)}{\|\mathbf{r}\|^{2}}$ (22)
-where $[\cdot]_{\times}$ denotes the skew-symmetric matrix representation:
+where $\cdot$$_{\times}$ is the skew-symmetric matrix representation:
 $[\mathbf{r}]_{\times}=\begin{bmatrix}0&-r_{3}&r_{2}\\ r_{3}&0&-r_{1}\\ -r_{2}&r_{1}&0\end{bmatrix}$ (23)
-where $r_{i}$ is the $i$-th element of $\mathbf{r}$.
-3D rigid transformation is a special case of 3D affine transformation, whereas our method is capable of handling 3D affine transformation.
-Therefore, we aim to adapt our method from Section 5 to accommodate 3D rigid transformation as well.
-Here’s the procedure: We continue to use the algorithm outlined in Section 5 for registration, where the transformation is defined as a 3D affine transformation: $T(\mathbf{x}|\mathbf{R},\mathbf{t})=\mathbf{R}\mathbf{x}+\mathbf{t}$, with $\mathbf{R}$ representing the linear part of the affine transformation.
-The main differences from the algorithm in Section 5 are as follows: (1) Instead of branching over $\mathbf{R}$ and $\mathbf{t}$, we branch over $\mathbf{r}$ and $\mathbf{t}$.
-(2) In each iteration of the BnB algorithm, given the range of $\mathbf{r}$, we calculate the range of $\mathbf{R}$ from the range of $\mathbf{r}$ based on equation (21).
-(The method for efficiently computing the range of $\mathbf{R}$ is discussed in Section 7.1.)
-In the above algorithm, (20) is employed to compute the upper bound, assuming that the transformation can be affine.
-While methods exist (see Sec.
-5.3 of [6]) for computing the upper bound with a rigid transformation constraint, they often result in higher upper bounds, which slows down our method’s convergence.
-Hence, we opt to use (20) for computing the upper bound in this work.
-7.1 The range of $\mathbf{R}$ from the range of $\mathbf{r}$
-Given a range $[\underline{\mathbf{r}},\overline{\mathbf{r}}]$ for $\mathbf{r}$, the computation of the range of $\mathbf{R}$, as derived from (21), involves determining $R_{ij}$ values within the bounds specified by $\underline{\mathbf{r}}$ and $\overline{\mathbf{r}}$.
-This computation can be facilitated through optimization solvers such as the Matlab function $fmincon$.
-However, this process can be burdensome, particularly when executed iteratively as a subroutine of the Branch and Bound (BnB) algorithm.
-In order to alleviate this computational burden, we propose the utilization of precomputation techniques to enhance efficiency.
-Specifically, prior to initiating the BnB algorithm, we construct a regular grid, with a predefined width (set as $150$ in this study), over the initial cube $[\underline{\mathbf{r}}_{0},\overline{\mathbf{r}}_{0}]$.
-Subsequently, we compute the corresponding $R_{ij}$ values for all grid points and store them.
-During the execution of the BnB algorithm, when presented with a range $[\underline{\mathbf{r}},\overline{\mathbf{r}}]$, we first identify the grid points falling within this range.
-Subsequently, we determine the minimum and maximum precomputed $R_{ij}$ values for these points, effectively approximating $\min\{R_{ij}|\underline{\mathbf{r}}\leq\mathbf{r}\leq\overline{\mathbf{r}}\}$ and $\max\{R_{ij}|\underline{\mathbf{r}}\leq\mathbf{r}\leq\overline{\mathbf{r}}\}$.
+with $r_{i}$ being the $i$-th element of $\mathbf{r}$.
+We adapt our core algorithm from Section 5 to handle this non-linear rigid transformation. The key modifications are:
+  1. 1.
+    Instead of branching over the full affine parameters, we branch over the 6-dimensional rigid parameters: the angle-axis vector $\mathbf{r}$ and the translation vector $\mathbf{t}$.
+  2. 2.
+    In each iteration of the BnB algorithm, the range of the rotation matrix $\mathbf{R}$ is computed from the current range of the angle-axis vector $\mathbf{r}$ using the formula in Eq. (22).
+For the upper bound calculation, we continue to use the affine-based approach from Eq.
+(19).
+While a dedicated method for rigid transformations exists (see Sec.
+5.3 of [5]), it often yields higher upper bounds, which hinders the pruning process and slows down our algorithm’s convergence.
+Therefore, we prioritize speed by using the affine upper bound, even though it doesn’t strictly enforce the rigid constraint.
+7.1 Efficiently Computing the Range of $\mathbf{R}$
+Given a range $\underline{\mathbf{r}},\overline{\mathbf{r}}]$ for the angle-axis vector $\mathbf{r}$, finding the exact range of each element of the rotation matrix $\mathbf{R}$ is a computationally expensive non-linear optimization problem.
+Performing this calculation at every step of the BnB algorithm would be a significant bottleneck.
+To address this, we use a precomputation and grid-based approximation method.
+Before the BnB algorithm begins, we create a regular grid over the initial search space of $\mathbf{r}$.
+For each grid point, we compute the corresponding rotation matrix $\mathbf{R}$ and store the values.
+Then, during the BnB process, when we need to find the range of $\mathbf{R}$ for a given hypercube $\underline{\mathbf{r}},\overline{\mathbf{r}}]$, we simply find all precomputed grid points that fall within this hypercube and take the minimum and maximum of their corresponding $\mathbf{R}$ values.
+This provides a fast, though approximate, way to get the necessary bounds, significantly speeding up the overall algorithm.
 8 Experiments
-We implement the proposed algorithm, referred to as RPM-BP, using Matlab R2023b.
-To evaluate its performance, we conduct comparisons with other methods on a computer equipped with 6-core CPUs running at 3.2 GHz.
-For the competing methods that solely produce point correspondences, we utilize the generated point correspondences to compute the optimal affine transformation between two sets of points.
-The matching error is then quantified as the root-mean-square distance between the transformed model inliers and their corresponding scene inliers.
-To ensure efficient computation, we set the maximum branching depth of RPM-BP to 12.
+We implemented the proposed algorithm, which we refer to as RPM-BP, using Matlab R2023b. All comparative experiments were conducted on a computer equipped with 6-core CPUs running at 3.2 GHz.
+For competing methods that output only point correspondences, we used these correspondences to compute the optimal affine transformation between the two point sets.
+We defined the matching error as the root-mean-square distance between the transformed model inliers and their corresponding scene inliers.
+To ensure efficient computation, the maximum branching depth of RPM-BP was set to 12.
 8.1 2D Registration
-We conduct comparisons with RPM-HTB [6], RPM-PA [5] and RPM-CAV [4], all of which utilize global optimization techniques.
-These methods are adept at handling partial overlap and permit arbitrary similarity transformations between two point sets, rendering them suitable candidates for our comparative analysis.
+We compared our method with RPM-HTB [5], RPM-PA [4], and RPM-CAV [3].
+These methods were chosen because they are all based on global optimization techniques, can handle partial overlap, and allow for arbitrary similarity transformations, making them suitable for direct comparison with our approach.
 8.1.1 2D Synthetic Data
-Synthetic data offer a convenient means to assess specific facets of an algorithm’s performance.
-We conduct five distinct types of tests to gauge a method’s resilience to various disturbances: i) Deformation test: The prototype shape undergoes non-rigid deformation to generate the scene set.
-ii) Noise test: Positional noise is introduced to perturb the prototype shape, resulting in the scene point set.
-iii) Mixed outliers and inliers test: Random outliers are superimposed on the prototype shape to create both point sets.
-iv) Separate outliers and inliers test: Random outliers are added to different areas of the prototype shape, generating the two point sets.
-v) Occlusion+outlier test: Initially, the prototype shape is occluded to produce the two point sets.
-Subsequently, random outliers (with a fixed outlier-to-data ratio of 0.5) are incorporated into different regions of the point sets.
+Synthetic data allows us to systematically evaluate an algorithm’s resilience to specific disturbances. We performed five distinct types of tests:
+  1. 1.
+    Deformation test: The prototype shape is non-rigidly deformed to create the scene point set.
+  2. 2.
+    Noise test: Positional noise is added to the prototype shape to generate the scene point set.
+  3. 3.
+    Mixed outliers test: Random outliers are superimposed on both the prototype and scene shapes.
+  4. 4.
+    Separate outliers test: Random outliers are added to different areas of the prototype shape to generate the two point sets.
+  5. 5.
+    Occlusion + Outlier test: The prototype shape is occluded to produce two point sets, to which random outliers (with a fixed outlier-to-data ratio of 0.5) are then added in different regions.
 These tests are visually illustrated in Fig.
 7.
-Additionally, to assess a method’s capability in handling arbitrary rotation and scaling, random rotations and scalings within the range of $[0.5,1.5]$ are applied when generating the two point sets.
-Fig.
-7 provides examples of registrations performed by different methods.
-The registration errors produced by various methods are depicted in the top two rows of Fig.
+To assess a method’s ability to handle arbitrary rotations and scaling, random rotations and scalings within the range of $[0.5,1.5]$ were applied when generating the point sets.
+Examples of registration results from different methods are provided in Fig.
 7.
-The result suggests that, in comparison to alternative methods, RPM-BP and RPM-HTB, with $n_{p}$ chosen as the ground truth, exhibits robustness against deformation, noise, and outliers when the outliers are distinct from the inliers.
-However, their robustness diminishes when outliers are intermixed with inliers.
-This limitation arises due to the fact that our method lacks $\epsilon-$global optimality, as discussed in Sec.
-5.7, thereby affecting its robustness.
-Regarding different transformation options, RPM-BP, when employing similarity transformation, demonstrates greater resilience to various disturbances than when utilizing affine transformation.
-Furthermore, the choice of $n_{p}$ value significantly influences the performance of RPM-BP.
-Specifically, RPM-BP with an $n_{p}$ value close to the ground truth yields superior results.
-Moreover, RPM-BP displays less sensitivity to variations in the $n_{p}$ value compared to RPM-HTB, particularly evident in tests involving outliers and occlusion.
-The average run times for various methods are depicted in the bottom row of Fig.
+The registration errors produced by each method are shown in the top two rows of Fig.
 7.
-RPM-BP exhibits the best performance in scenarios involving deformation, positional noise, and mixed inlier-outlier tests.
-Comparing different transformation options, RPM-BP employing similarity transformation proves to be more efficient than when employing affine transformation.
-This efficiency is attributed to the lesser number of parameters involved in similarity transformation.
-The outlier tests also provide insights into how various methods scale with problem size. Among them, RPM-BP, RPM-HTB, and RPM-CAV demonstrate the best scalability, followed by RPM-PA.
-Refer to caption Refer to caption Refer to caption
+Our results indicate that RPM-BP and RPM-HTB, when the number of correspondences $n_{p}$ is chosen as the ground truth, are robust against deformation, noise, and separate outliers.
+However, their performance degrades when outliers are mixed with inliers.
+This limitation is a consequence of our method not being $\epsilon$-globally optimal, as discussed in Sec.
+5.7.
+When using different transformation types, RPM-BP demonstrates greater resilience to disturbances when employing a similarity transformation compared to an affine transformation.
+Furthermore, the performance of RPM-BP is significantly influenced by the chosen $n_{p}$ value, with results improving as $n_{p}$ approaches the ground truth.
+We also found that RPM-BP is less sensitive to variations in $n_{p}$ than RPM-HTB, especially in outlier and occlusion tests.
+The average run times are depicted in the bottom row of Fig.
+7.
+RPM-BP is the most efficient method in the deformation, positional noise, and mixed inlier-outlier tests.
+Comparing transformation types, RPM-BP is more efficient with similarity transformations due to the smaller number of parameters.
+Finally, the outlier tests also provide insight into how each method scales with problem size. RPM-BP, RPM-HTB, and RPM-CAV demonstrate the best scalability, followed by RPM-PA.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 Figure 5: a) to (c): Model point sets and examples of scene point sets in the deformation and noise tests, respectively.
@@
 In all cases, model points are indicated by red circles, while scene points are represented by blue crosses.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 6: Average registration errors (top 2 rows) and run times (bottom row) by various methods under various $n_{p}$ values (ranging from $1/2$ to $1/1$ of the ground truth value) over 100 random trials for 2D deformation, positional noise, mixed outliers and inliers, separate outliers and inliers, and occlusion+outlier tests.
+(a)RPM-BP(simi) (b)RPM-BP(aff) (c)RPM-HTB(simi) (d)RPM-HTB(aff) (e)RPM-PA (f)RPM-CAV
+Figure 7: Example of registration results from different methods in the separate outliers and inliers test, with $n_{p}$ chosen as ground truth for all methods.
+8.1.2 2D Real-World Data
+Point sets extracted from images represent a realistic and challenging data type for evaluating the robustness and generalizability of registration methods, as they inherently contain noise, irregular contours, and complex structural features.
+Experimental Setup
+To generate these point sets, we utilized the standard Canny edge detector on selected images sourced from the Caltech-256 [32] and VOC2007 [33] datasets.
+A critical test of rotational invariance was implemented by subjecting the model point set to a deliberate $180^{\circ}$ rotation prior to registration against the scene point set.
+For consistency, a Similarity transformation model was assumed and used for all tested methods.
+The visual setup for this experiment is detailed in Fig.
+9, with final alignment results shown in Fig.
+9.
+Results
+The registration outcomes presented in Fig.
+9 demonstrate that our method consistently achieves a tighter alignment to the underlying target objects than competing algorithms.
+This advantage is notably pronounced in the bike image test.
+Figure 8: Model images with model point sets superimposed.
 Refer to caption Refer to caption
 Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 6: Average registration errors (top 2 rows) and run times (bottom row) by RPM-BP, RPM-HTB [6], and RPM-CAV [4] under various $n_{p}$ values (ranging from $1/2$ to $/1$ 1 of the ground truth value) over 100 random trials for 2D deformation, noise, mixed outliers and inliers, separate outliers and inliers, and occlusion+outlier tests.
-(a) RPM-BP (simi) (b) RPM-BP (aff) (c) RPM-HTB (simi) (d) RPM-HTB (aff) (e) RPM-PA (f) RPM-CAV
-Figure 7: Example of registration results from different methods in the separate outliers and inliers test, with $n_{p}$ chosen as ground truth for all methods.
+Figure 9: For each category: scene images with scene point sets superimposed, registration results by RPM-BP, RPM-HTB [5], RPM-PA [4] and RPM-CAV [3] using similarity transformation.
+The $n_{p}$ value for each method is chosen as $0.9$ the minimum of the cardinalities of two point sets.
 8.2 3D Registration
 8.2.1 3D Synthetic Data
-Similar to the experimental setup detailed in Sec.
-8.1.1, we conducted five types of tests to assess a method’s resilience against various disturbances.
-These tests include: i) Deformation test, ii) Noise test, iii) Mixed outliers and inliers test, iv) Separate outliers and inliers test, and v) Occlusion+outlier test.
-Refer to Fig.
-8 for an illustration.
-Additionally, examples of registration by different methods are depicted in Fig.
-10.
-The top 2 rows of Fig.
-9 displays registration errors from various methods.
-Results show that RPM-BP and RPM-HTB, with $n_{p}$ chosen as the ground truth, exhibits robustness against deformation, noise, and outliers when outliers are separate from inliers.
-However, they loses robustness when outliers are mixed with inliers.
-In contrast, GORE and TEASER++ struggle with handling deformation, noise, and outliers, especially when outliers are mixed with inliers, likely due to their reliance on features.
-In terms of different choices of $n_{p}$, RPM-BP performs significantly better when $n_{p}$ closely matches the ground truth.
+Similar to the 2D experiments, we conducted five types of tests to evaluate each method’s resilience to various disturbances: i) Deformation, ii) Noise, iii) Mixed outliers, iv) Separate outliers, and v) Occlusion + Outlier tests.
+Fig.
+12 provides a visual illustration of these tests, while Fig.
+12 shows examples of registration results from different methods.
+The registration errors for various methods are shown in the top two rows of Fig.
+12.
+The results indicate that RPM-BP and RPM-HTB, when $n_{p}$ is set to the ground truth value, are robust against deformation, noise, and separate outliers.
+However, their robustness diminishes when outliers are mixed with inliers.
+In contrast, GORE and TEASER++ struggle with deformation, noise, and mixed outliers, likely because they rely on features.
+As in the 2D case, the performance of RPM-BP improves significantly when $n_{p}$ closely matches the ground truth.
 The bottom row of Fig.
-9 displays the average run times for various methods.
-TEASER++ emerges as the fastest, with Go-ICP following closely behind, and RPM-BP and RPM-HTB trailing after.
-GORE consistently exhibits slower performance across most cases.
-Notably, FRS experiences a rapid decline in efficiency when faced with challenging registration problems, such as an increase in outliers in both types of outlier tests.
-Examining the two types of outlier tests also reveals how different methods scale with problem size.
-Go-ICP and GORE demonstrate the best scalability with problem size, followed by TEASER++, RPM-BP and RPM-HTB.
-Conversely, FRS exhibits the poorest scalability with problem size.
-Refer to caption Refer to caption
-Figure 8: (a) to (c): Model point sets and corresponding examples of scene point sets in the deformation and noise tests, respectively.
+12 shows the average run times.
+TEASER++ is the fastest, followed by Go-ICP, and then RPM-BP and RPM-HTB.
+GORE is consistently slower in most cases, while FRS experiences a rapid decrease in efficiency when the registration problem becomes more challenging, such as with an increasing number of outliers.
+In terms of scalability with problem size, Go-ICP and GORE perform best, followed by TEASER++, RPM-BP, and RPM-HTB. FRS, on the other hand, shows the poorest scalability.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 10: (a) to (c): Model point sets and examples of scene point sets in the deformation and noise tests, respectively.
 (d) to (i): Examples of model and scene point sets in the mixed outliers and inliers test ((d), (e)), separate outliers and inliers test ((f), (g)), and occlusion+outlier test ((h), (i)), respectively.
 In all cases, model points are indicated by red circles, while scene points are represented by blue crosses.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 9: Average registration errors (top 2 rows) and run times (bottom row) of RPM-BP, RPM-HTB, Go-ICP, FRS, GORE, and TEASER++ under varying $n_{p}$ values (ranging from $1/2$ to $/1$ 1 of the ground truth value) across 100 random trials for 3D deformation, noise, mixed outliers and inliers, separate outliers and inliers, and occlusion+outlier tests.
-For the separate outliers and inliers and occlusion+outlier tests, only a portion of FRS’s run time results is presented, as FRS experiences significant slowdowns with large problem sizes.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 11: Average registration errors (top 2 rows) and run times (bottom row) of various methods under varying $n_{p}$ values ($1/2$ or $1/1$ of the ground truth value) across 100 random trials for 3D deformation, positional noise, mixed outliers and inliers, separate outliers and inliers, and occlusion+outlier tests.
       (a) RPM-BP (b) RPM-HTB (c) Go-ICP (d) FRS (e) GORE (f) TEASER++
-Figure 10: Examples of registration results from different methods in the separate outliers and inliers test, where the $n_{p}$ values of RPM-HTB and Go-ICP are both chosen to match the ground truth.
+Figure 12: Examples of registration results from different methods in the separate outliers and inliers test, where the $n_{p}$ values of RPM-HTB and Go-ICP are both chosen as the ground truth.
 8.2.2 3DMatch Dataset
-The 3DMatch benchmark [41] encompasses scans of 62 scenes sourced from five established RGB-D reconstruction datasets (sun3d, 7-scenes, rgbd-scenes-v2, bundlefusion, and analysis-by-synthesis).
-D3Feat [42] offers a unified file structure and format.
-For convenience, we utilize the scan pairs provided by D3Feat to assess the performance of various methods.
-The matching errors from different methods are depicted in Fig.
-12.
-The results indicate that RPM-BP and RPM-HTB generally outperform Go-ICP and FRS but lag behind GORE and TEASER++.
-The superior performance of GORE and TEASER++ is partly attributed to their use of features, which significantly enhances their capability to match complex structures.
-In contrast, other methods rely solely on point position information.
-Additionally, the 3DMatch benchmark comprises purely rigid transformations, aligning well with the assumptions of GORE and TEASER++.
-Among RPM-BP and RPM-HTB, RPM-BP exhibits slightly superior performance, particularly evident in the sun3d and rgbd-scenes-v2 tests.
-Examples of registration by different methods are presented in Fig. 12.
-Refer to caption Refer to caption
+The 3DMatch benchmark [34] provides scans from 62 scenes across five established RGB-D reconstruction datasets (sun3d, 7-scenes, rgbd-scenes-v2, bundlefusion, and analysis-by-synthesis).
+For convenience, we used the scan pairs from D3Feat [35] to evaluate the performance of our method.
+Fig.
+14 shows the matching errors for the different methods.
+The results indicate that RPM-BP and RPM-HTB generally outperform Go-ICP and FRS but are less accurate than GORE and TEASER++.
+The superior performance of GORE and TEASER++ can be attributed to their use of features, which significantly enhances their ability to match complex structures.
+In contrast, the other methods rely solely on point position information.
+Furthermore, the 3DMatch benchmark contains purely rigid transformations, which aligns perfectly with the assumptions of GORE and TEASER++.
+Between our method and RPM-HTB, RPM-BP demonstrates slightly better performance, especially in the sun3d and rgbd-scenes-v2 tests.
+Examples of registration results are presented in Fig. 14.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 Figure 13: Average registration errors from various methods across five RGB-D reconstruction datasets.
@@
 9 Conclusion
-In this paper, we introduced a BnB-based point set registration algorithm capable of aligning partially overlapping point sets while remaining invariant to the corresponding transformation.
-The method offers several advantages: the lower bound can be efficiently computed via linear assignment and low-dimensional convex quadratic program, and the dimensionality of the branching space matches the number of transformation parameters.
-These attributes make the method computationally efficient and scalable with problem size.
-By employing the BnB algorithm, our method demonstrates robustness against non-rigid deformations, positional noise, and outliers, particularly when outliers are distinct from inliers compared to state-of-the-art approaches.
-However, the proposed method is limited to transformations with few parameters due to the exponential complexity of the BnB algorithm with respect to the dimensionality of the branching space.
-Additionally, the method’s robustness is significantly affected by the parameter setting for $n_{p}$.
-In the future, we plan to develop an adaptive scheme for setting $n_{p}$.
-ACKNOWLEDGMENTS
-This work was supported by the Fundamental Research Program of Shanxi Province (202103021223381, 202303021222267 and 202303021222271) and Science and Technological Innovation Program of Higher Education Institutions in Shanxi under Grant (2023L335)
-Appendix A A weak result about positive semidefiniteness of matrix $\mathbf{H}^{0}+\mathbf{C}$
+In this paper, we presented RPM-BP, a novel Branch-and-Bound (BnB) algorithm for globally optimal point set registration under partial overlap.
+Our method is designed to be invariant to the underlying transformation, encompassing both 2D similarity/affine and 3D rigid transformations.
+The efficiency of RPM-BP stems from several key features:
+  1. 1.
+    Efficient Lower Bound Computation: The lower bound is efficiently computed by decoupling the minimization problem into a tractable linear assignment problem (for correspondences) and a low-dimensional convex quadratic program (for transformation parameters).
+  2. 2.
+    Low-Dimensional Branching Space: The dimensionality of the branching space is confined solely to the transformation parameters, which is inherently low-dimensional (e.g., 4 for 2D similarity, 6 for 3D rigid).
+This greatly mitigates the exponential complexity typical of BnB algorithms.
+These attributes contribute to the method’s computational efficiency and its strong scalability with respect to the number of points.
+Our extensive experiments confirm that RPM-BP demonstrates superior robustness against non-rigid deformations, positional noise, and outliers, particularly in scenarios where outliers and inliers are spatially separated.
+Despite its advantages, the proposed method has two primary limitations.
+First, its use is inherently restricted to transformations with a small number of parameters, as the complexity of the BnB algorithm remains exponential with respect to the branching space dimensionality.
+Second, the method’s overall robustness is sensitive to the parameter setting for the number of inliers, $n_{p}$.
+For future work, we plan to focus on developing an adaptive scheme for setting the parameter $n_{p}$ to enhance the method’s performance and practical applicability.
+Appendix A Heuristic Justification for the Positive Semidefiniteness of $\mathbf{H}^{0}+\mathbf{C}$
+The convergence of our Branch-and-Bound algorithm relies on the lower bound subproblem being convex, which requires the matrix $\mathbf{H}^{0}+\mathbf{C}$ to be Positive Semidefinite (PSD).
+While a formal proof that $\mathbf{H}^{0}+\mathbf{C}\succeq 0$ is elusive, we offer a heuristic justification, consistent with our empirical findings in Section 5.6.
 Proposition 1.
-Matrix $\mathbf{H}^{0}+\mathbf{C}$ is close to being positive semidefinite.
+The matrix $\mathbf{H}^{0}+\mathbf{C}$ is an average of matrices that are heuristically expected to be close to Positive Semidefinite.
 Proof.
-For any $\mathbf{p}\in\Omega$, it is apparently that $\mathbf{p}\geq 0$.
-Thus, given $\mathbf{p}\in\Omega$, the function $E$ in equation (3a) is a convex quadratic function of $\boldsymbol{\theta}$.
-Therefore, from equation (5), we can infer that $\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})+\mathbf{C}\succeq 0$ for any $\mathbf{p}\in\Omega$.
-Let
-$\mathbf{p}^{ij}=\arg\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$
-Apparently ${\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p}^{ij})}+\mathbf{C}\succeq 0$.
-Based on the fact that $\underline{\Gamma}_{ij}=\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}=\min_{\mathbf{p}\in\{\mathbf{p}^{ij}\}}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$, we have that $\underline{\boldsymbol{\Gamma}}$ is close to one of $\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p}^{ij})$, $\forall i,j$.
-Consequently $\underline{\boldsymbol{\Gamma}}+\mathbf{C}$ is close to being positive semidefinite.
-In an analogous way, let
-$\mathbf{q}^{ij}=\arg\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$
-Apparently ${\text{mat}(\mathbf{K}\mathbf{B}_{2}q^{ij})}+\mathbf{C}\succeq 0$.
-Based on $\overline{\Gamma}_{ij}=\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}=\max_{\mathbf{p}\in\{\mathbf{q}^{ij}\}}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$, we have $\underline{\boldsymbol{\Gamma}}$ is close to one of $\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{q}^{ij})$, $\forall i,j$.
-Consequently $\overline{\boldsymbol{\Gamma}}+\mathbf{C}$ is close to being positive semidefinite.
-Therefore $\mathbf{H}^{0}+\mathbf{C}=\frac{\left(\underline{\boldsymbol{\Gamma}}+\mathbf{C}\right)+\left(\overline{\boldsymbol{\Gamma}}+\mathbf{C}\right)}{2}$ is close to being positive semidefinite.
+We first recall the definition of the Hessian matrix $\mathbf{G}(\mathbf{p})$ with respect to $\boldsymbol{\theta}$ for a fixed correspondence vector $\mathbf{p}$:
+$\mathbf{G}(\mathbf{p})=\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})+\mathbf{C}.$
+Since the point correspondence vector $\mathbf{p}$ satisfies $\mathbf{p}\geq 0$ (i.e., $\mathbf{p}\in\Omega$), and the original objective function $E(\mathbf{p},\boldsymbol{\theta})$ represents a weighted sum of squared errors, it is fundamentally a convex quadratic function of $\boldsymbol{\theta}$ for any feasible $\mathbf{p}$.
+This mandates that the Hessian must be Positive Semidefinite for all valid $\mathbf{p}$:
+$\mathbf{G}(\mathbf{p})\succeq 0\quad\text{for all }\mathbf{p}\in\Omega.$
+The matrix $\mathbf{H}^{0}+\mathbf{C}$ is defined as the average of two bounding matrices:
+$\mathbf{H}^{0}+\mathbf{C}=\frac{1}{2}\left[(\underline{\boldsymbol{\Gamma}}+\mathbf{C})+(\overline{\boldsymbol{\Gamma}}+\mathbf{C})\right].$
+where $\underline{\boldsymbol{\Gamma}}+\mathbf{C}$ (and similarly $\overline{\boldsymbol{\Gamma}}+\mathbf{C}$) is an element-wise lower bound on the Hessian $\mathbf{G}(\mathbf{p})$:
+$(\underline{\boldsymbol{\Gamma}}+\mathbf{C})_{ij}=\left(\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}\right)+\mathbf{C}_{ij}.$
+The element-wise minimum $\underline{\boldsymbol{\Gamma}}_{ij}$ (and maximum $\overline{\boldsymbol{\Gamma}}_{ij}$) is attained by some $\mathbf{p}\in\Omega$.
+Since all $\mathbf{G}(\mathbf{p})$ are PSD matrices, the elements used to construct $\underline{\boldsymbol{\Gamma}}+\mathbf{C}$ and $\overline{\boldsymbol{\Gamma}}+\mathbf{C}$ are drawn from a collection of PSD matrices.
+While $\underline{\boldsymbol{\Gamma}}+\mathbf{C}$ and $\overline{\boldsymbol{\Gamma}}+\mathbf{C}$ may not be strictly PSD themselves (as the minimum/maximum of each entry may be attained by a different $\mathbf{p}$), they are structurally derived from PSD matrices.
+It is therefore reasonable to conclude that both $(\underline{\boldsymbol{\Gamma}}+\mathbf{C})$ and $(\overline{\boldsymbol{\Gamma}}+\mathbf{C})$ are close to being positive semidefinite.
+Because the sum of two PSD matrices is PSD, their average, $\mathbf{H}^{0}+\mathbf{C}$, is consequently expected to be close to being positive semidefinite, ensuring the practical convexity of the subproblem.
+∎